### v1

# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. Their potential in aiding quantum software development remains underexplored, particularly for the PennyLane framework—a leading platform for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific code samples of quantum circuits and their contextual descriptions, specifically curated to train/fine-tune LLM-based quantum code assistance. Our key contributions are threefold: (1) the automatic creation and open-source release of a comprehensive PennyLane dataset leveraging quantum computing textbooks, official documentation, and open-source repositories; (2) the development of a systematic methodology for data refinement, annotation, and formatting to optimize LLM training efficiency; and (3) a thorough evaluation, based on a Retrieval-Augmented Generation (RAG) framework, demonstrating the effectiveness of our dataset in streamlining PennyLane code generation and improving quantum development workflows. Compared to existing efforts that predominantly focus on Qiskit, our dataset significantly broadens the spectrum of quantum frameworks covered in AI-driven code assistance. By bridging this gap and providing reproducible dataset-creation methodologies, we aim to advance the field of AI-assisted quantum programming, making quantum computing more accessible to both newcomers and experienced developers.
Index Terms:
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), Code Completion and Assistance, Dataset Curation and Annotation, Evaluation Framework
I Introduction
Quantum computing has emerged as a transformative technology, with the potential to solve complex problems that are infeasible for classical systems [1, 2]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [3]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [4, 5, 6, 7, 8].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks, hindering widespread adoption [9, 10]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [11]. However, unlike the Qiskit framework [12], which benefits from dedicated AI-driven code assistants [13], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has witnessed a surge in the capabilities of LLMs, which excel at reasoning, contextual understanding, and automated code generation [14, 15, 16]. These models have been successfully leveraged in classical programming [17, 18], improving developer efficiency and reducing syntactical and logical errors. However, their application in quantum programming remains largely underexplored, particularly for PennyLane. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for its refinement, annotation, formatting, and evaluation to facilitate LLM training.
State-of-the-Art and Limitations: While several studies have explored LLM-driven code generation for classical programming tasks [19, 20], research on quantum-specific datasets remains scarce. The most notable effort in this domain is the Qiskit Code Assistant [13], which leverages a Qiskit-specific dataset for AI-assisted quantum code generation. However, its hardware-centric approach does not extend to PennyLane, which specializes in variational quantum circuits and quantum machine learning applications [21].
The primary challenge in developing a PennyLane code assistant lies in data availability. Unlike classical programming languages, where vast repositories of well-documented code exist, quantum computing resources are fragmented across research papers, user forums, open-source repositories, and official documentation [22]. Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
Our Approach and Contributions: In this work, we introduce PennyLang to address the gap in LLM-assisted quantum code generation for pennylane through the following novel contributions. An overview of our contributions is presented in Fig. 1.
Figure 1: Overview of novel contributions in this work.
  1. 1.
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, two quantum computing textbooks, and the official PennyLane documentation. Each sample is meticulously annotated with explanatory comments and contextual descriptions, facilitating more effective LLM training for quantum programming.
  2. 2.
    Systematic Data Refinement and Annotation: We develop a robust methodology for dataset cleaning, eliminating non-PennyLane snippets and enriching the samples with structured metadata. To further enhance usability, we employ the GPT-4o API to convert code samples into an instruction-query format, providing clear guidance for model training in a structured, task-oriented manner.
  3. 3.
    RAG-Based Evaluation Framework: We use a Retrieval-Augmented Generation (RAG) evaluation framework to assess the dataset’s impact on model performance. The framework compares outputs from three models—GPT-4o mini, Claude 3.5, and Qwen 7B—under two conditions: (i) with access to the dataset as retrieval-augmented context and (ii) without dataset augmentation. The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
Summary of Key Results: RAG-enhanced quantum code generation significantly improved performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability.
Paper Organization: Sec. II provides background information and reviews related work. Sec. III outlines our methodology for dataset collection, cleaning, and annotation along with the tokenization and formatting pipeline used for preprocessing. Sec. IV presents experimental results, including an evaluation of our dataset using a RAG-based framework. Sec. V presents some case studies Finally, Sec. VI concludes the paper with a summary of findings and future research directions.
II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [23] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [24] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [25] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [26] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [27] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [28] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [29] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [30] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [31] and automated code generation [32], their application in quantum software development remains under-explored. Existing studies on AI-driven code assistants primarily focus on classical programming languages such as Python, Java, and C++, or quantum computing frameworks like Qiskit [13]. However, the PennyLane framework [11], widely used for hybrid quantum-classical computing, lacks dedicated research efforts in AI-assisted quantum programming.
This section presents an in-depth review of related work in three key areas:
  1. 1.
    LLMs for Code Generation in Classical Programming
  2. 2.
    LLMs for Quantum Programming
  3. 3.
    Datasets for Training LLMs in Quantum Code Generation (summarized in Table I)
We then summarize existing contributions and highlight our novel approach in developing a PennyLane-centric dataset to advance AI-driven quantum programming, as shown in Table II.
II-A LLMs for Code Generation in Classical Programming
Transformer-based models have significantly impacted AI-driven programming, automating code generation, completion, and debugging. Key contributions include:
  • Codex (OpenAI) [17] introduced a language model trained on public code repositories, enabling natural language-to-code translation.
  • CodeBERT [33] leveraged bidirectional transformers to enhance source code understanding.
  • InCoder [20] developed a multi-turn interactive model for code infilling and synthesis.
  • AlphaCode [16] showed competitive programming-level LLM capabilities.
Additionally, benchmarking studies have evaluated LLM performance on programming tasks:
  • HumanEval [15] proposed an execution-based evaluation for Python code generation.
  • MBPP [34] introduced a multi-language benchmark covering correctness and execution constraints.
  • CodeXGLUE [35] aggregated multiple datasets to benchmark code-related tasks.
These studies highlight the necessity of well-structured datasets to enhance LLM code generation performance.
II-B LLMs for Quantum Programming
Existing research on AI-powered quantum code assistants focuses primarily on the Qiskit framework:
  • Qiskit Code Assistant (IBM Research) [13] assists users in debugging and optimizing Qiskit code.
  • Qiuchi et al. [36] adapted BERT-based models to improve quantum code understanding.
  • CodeLlama-Quantum (Meta AI) [18] extended CodeLlama for Qiskit-based applications.
Despite these advancements, challenges persist:
  1. 1.
    Limited Training Data: Quantum programming lacks large-scale, high-quality datasets.
  2. 2.
    Complexity of Quantum Circuits: Unlike classical code, quantum programs require understanding quantum mechanics.
  3. 3.
    Hybrid Workflows: PennyLane integrates classical and quantum processing, requiring LLMs to learn hybrid paradigms.
These challenges underscore the need for PennyLane-specific datasets for LLM training, as none currently exist, reinforcing the importance of our research.
Figure 2: Methodology for PennyLang quantum code generation and evaluation. Data is collected from GitHub, quantum books, and PennyLane docs, then refined, tokenized, and formatted for LLMs. The curated dataset is processed by GPT-4o for input-output conversion. User queries are handled via a retrieval-augmented generation (RAG) framework using LangChain and Chroma to fetch relevant code. Responses from GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B-Instruct-Turbo are compared against non-RAG outputs and evaluated on functionality, syntax, and modularity to assess RAG’s impact.
Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
TABLE II: Comparison of Related Work and Our Contributions
Research Area Existing Works Our Contributions
LLMs for Code Generation Codex [17], CodeBERT [33] PennyLane-specific dataset
LLMs for Quantum Programming Qiskit Assistant [13], Qiuchi et al. [36] Hybrid quantum-classical dataset
Quantum Code Datasets QCircuitNet [28], QDataset  [28] First PennyLane dataset
III PennyLang Framework
III-A Methodology for Data Collection and Refinement
Fig. 2 illustrates the methodology employed to construct a high-quality dataset of PennyLane code samples, ensuring that each snippet is genuinely relevant, well-structured, and accurately annotated. Our data collection strategy focused on extracting, filtering, and refining code snippets from multiple sources, including GitHub repositories, quantum computing books, and official PennyLane documentation.
III-A1 Data Collection from GitHub Repositories
To systematically gather PennyLane-specific code, we utilized the GitHub API to search for publicly accessible repositories with permissive open-source licenses (e.g., MIT, Apache 2.0, BSD). The selection criteria required that the term "PennyLane" be present in either the repository name, description, or README file.
To maintain dataset quality and avoid redundancy, we followed these constraints:
  • Only repositories from the main branch at the latest commit were considered.
  • Forked repositories were systematically excluded.
  • Only .py files (Python scripts) were extracted for further analysis.
  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
  • Large, autogenerated files (e.g., test logs, package files) were excluded.
Since PennyLane maintains backward compatibility and rarely deprecates features [11], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code.
This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [30].
III-A2 Extracting Quantum Code from Books
In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books:
  • “Quantum Machine Learning: An Applied Approach” [21]
  • “A Practical Guide to Quantum Machine Learning and Quantum Optimization” [37]
These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
  1. 1.
    Identifying all PennyLane-related code snippets.
  2. 2.
    Diversity Assurance: The dataset includes a broad spectrum of PennyLane use cases, from basic gate operations to advanced quantum optimization problems.
  3. 3.
    Extracting only the relevant code while discarding unrelated sections.
  4. 4.
    Manually verification of each snippet with explanations from the books, converting descriptive text into structured comments.
Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code). Thus, manual review was necessary to ensure that annotations correctly corresponded to each snippet.
III-A3 Scraping Code from PennyLane Documentation
The official PennyLane documentation [29] serves as a comprehensive reference for quantum programming, covering:
  • Quantum circuits and variational algorithms.
  • Quantum operators and templates.
  • Hybrid quantum-classical optimization.
  • Quantum chemistry applications.
TABLE III: Tokenized Input and Attention Masking
Tokenized Input Attention Mask
"<pad>" 0
"<pad>" 0
"def" 1
":" 1
"circuit" 1
Tokenized Input Attention Mask
"(" 1
")" 1
"return" 1
"qml.expval" 1
To further enhance our dataset, we systematically scraped all code snippets from the official PennyLane website. This included examples embedded in tutorials and API references.
As with book sources, documentation explanations were manually converted into comments, ensuring that each code sample contained self-contained, contextualized information. This practice enhances LLM interpretability, as models trained on such data can generate both code and explanations.
III-A4 Dataset Summary and Statistics
The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
TABLE IV: Dataset Composition
Source Samples Collected
Unofficial GitHub repositories 1,952
GitHub (PennyLaneAI repository + others) 1,321
Official PennyLane Documentation 53
Quantum Computing Books 21
Total Dataset Size 3,347
Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for fine-tuning large language models (LLMs) for PennyLane-based quantum programming.
III-A5 Ensuring Data Integrity and Diversity
To maintain dataset integrity, we applied multiple quality control measures:
  1. 1.
    Duplicate Removal: Identical code samples from different sources were eliminated.
  2. 2.
    Code Formatting: Samples were formatted consistently to follow PEP 8 style guidelines [38].
  3. 3.
    Manual Review: A human-in-the-loop verification was conducted to assess annotation accuracy.
These efforts ensure that the dataset is rich, well-structured, and optimized for AI model training.
III-A6 Implications for AI-Driven PennyLane Code Generation
The collected dataset significantly enhances AI-driven PennyLane code assistance in multiple ways:
  • Improving Model Performance: High-quality annotations ensure that models understand both syntax and context.
  • Facilitating Explainable AI: Models trained on our dataset can generate explanatory comments alongside code.
This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
Component Function Description
DirectoryLoader Document Loader Loads PennyLane dataset samples as text files from a directory.
Chroma Vector Database Stores embeddings for retrieval using similarity search.
OpenAIEmbeddings Embedding Model Converts text into vector embeddings for indexing and retrieval.
ChatPromptTemplate Prompt Formatting Defines structured prompts for AI models to ensure consistency.
create_retrieval_chain Retrieval Pipeline Creates a pipeline that integrates stored embeddings for improved responses.
ChatOpenAI OpenAI LLM Interface Connects to OpenAI models (GPT-4o-mini) for response generation.
ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework
Model Parameters Architecture Provider Usage in Pipeline
Qwen2.5-7B-Instruct-Turbo 7B Causal Decoder-Only Transformer Alibaba Cloud Generates PennyLane code
GPT-4o Mini Not disclosed Causal Decoder-Only Transformer OpenAI Evaluates and scores generated responses
Claude 3.5 Sonnet Not disclosed Causal Decoder-Only Transformer Anthropic Alternative model for response generation
III-B Tokenization and Formatting for LLM Training
This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Handling Variable-Length Code Samples
The collected PennyLane code samples exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
III-B2 Padding Strategies and Tokenizer Variations
Padding strategies differ across tokenizers:
  • Token-Specific Padding: Some tokenizers have a dedicated padding token, automatically appended to shorter sequences. For example, Hugging Face’s Tokenizer uses "<pad>" by default [39].
  • Manually Defined Padding: If a tokenizer lacks an explicit padding token, a placeholder token (e.g., "0" or "<mask>") must be manually assigned.
  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [40], ensuring that padding tokens do not interfere with causal attention mechanisms.
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, and Claude 3.5 Sonnet. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining autoregressive consistency.
III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [41] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table III represents the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
  • Attention Mask Structure: A binary mask is assigned to each token:
    • 1 → Meaningful tokens to be processed by the model.
    • 0 → Padding tokens ignored by the transformer.
  • Impact on Training: The self-attention mechanism [40] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
III-C Methodology for Dataset Evaluation
III-C1 Retrieval-Augmented Generation (RAG) Framework
To enhance LLM-based quantum code generation, we employ a Retrieval-Augmented Generation (RAG) methodology that integrates retrieval-based context with language model inference. This hybrid approach significantly improves:
  • Accuracy: By grounding responses in retrieved dataset samples, reducing model hallucination.
  • Adherence to PennyLane’s best practices: Ensuring generated quantum code aligns with official standards.
  • Response Quality: Structured retrieval enriches contextual understanding, improving coherence and correctness.
III-C2 RAG Implementation using LangChain
LangChain is utilized to construct a modular and efficient RAG pipeline. The framework consists of the following core components:
  1. 1.
    Dataset Loading: PennyLane dataset samples are extracted from curated dataset structured in intruction-response format.
  2. 2.
    Embedding Generation: Each sample is tokenized and converted into high-dimensional vector embeddings using OpenAIEmbeddings.
  3. 3.
    Vector Storage and Retrieval: The embeddings are stored in a Chroma vector database for fast retrieval.
  4. 4.
    Prompting and Response Generation:
    • When a query is submitted, Maximum Marginal Relevance (MMR) search fetches the most relevant samples in RAG-augmented input.
    • Retrieved context is formatted into a structured prompt.
    • The LLM model generates responses using both RAG-augmented and non-RAG prompts.
  5. 5.
    Evaluation: The responses are compared using a second LLM (GPT-4o Mini), which scores them based on:
    • Functionality: Logical correctness of the quantum circuit.
    • Syntax: Validity of the PennyLane code.
    • Modularity: Adherence to best practices in structuring quantum workflows.
Figure 3: Performance of GPT 4o mini, Claude, and Qwen With and Without RAG. Mean scores (±1 SD) over 29 trials for GPT 4o mini, Claude, and Qwen under both conditions (blue: With RAG; orange: Without RAG), with individual scores shown as swarmplot markers.
Figure 3 presents the results of our RAG-based evaluation methodology, demonstrating a significant performance improvement with retrieval-augmented responses. Each model was tested with 30 code samples, and responses were evaluated using GPT-4o, which graded them based on syntax, functionality, and modularity. The results indicate that the RAG-enhanced approach consistently outperforms the base LLM, yielding higher overall scores.
III-C3 LLM Models and Configuration
The models used in this framework differ in size, architecture, and inference capabilities. Table VI provides a comprehensive comparison of key characteristics including parameter count, context window size, and quantum-specific training data percentage.
Figure 4: Distribution of instruction and response lengths with their correlation. The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
III-C4 Prompt Design for RAG and Non-RAG Evaluation
To ensure fair evaluation, we construct two different prompts—one for RAG-augmented inference and another for direct model inference without retrieval.
  • Non-RAG Prompt:
    Base LLM Prompt (Without RAG) "You are an advanced assistant for quantum computing tasks using PennyLane. Provide a response to the user’s question. Make sure to follow best practices and optimize for efficiency. If you are unsure, state that clearly."
  • RAG-Enhanced Prompt:
    RAG-Enhanced Prompt with Context Retrieval "You are an advanced assistant for quantum computing tasks using PennyLane. Use the following pieces of retrieved context to answer the question. Use multiple retrieved code snippets whenever necessary to improve accuracy. If a single snippet does not provide enough information, retrieve and combine relevant parts from multiple sources. Make sure to follow best practices and optimize for efficiency. If you don’t know the answer, say that you don’t know." "{context}"
III-C5 Evaluation Criteria and Scoring
To assess the effectiveness of RAG-based retrieval, generated responses are evaluated using GPT-4o with the following scoring prompt:
Evaluation Prompt for GPT-4o You are an AI assistant tasked with evaluating two different answers to the same question. Question: {question} Answer with RAG: {rag_answer} Answer without RAG: {non_rag_answer} Please provide a score from 1 to 30 for each answer based on: - Functionality (correctness of quantum operations) - Syntax (validity of PennyLane code) - Modularity (code organization and efficiency) Justify your scoring.
III-C6 Findings and Impact of RAG Integration
After evaluating multiple quantum programming queries using the RAG vs. Non-RAG framework, we observe: Based on the evaluation results, RAG-enhanced responses consistently outperformed non-RAG responses across all tested models—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5. On average, RAG integration improved response scores by 11.67% in GPT-4o Mini, 7.69% in Claude 3.5, and 14.38% in Qwen 2.5. The improvement was most noticeable in functionality and syntax, as retrieved context helped prevent incorrect quantum operations and improved PennyLane-specific syntax accuracy. Claude 3.5 Sonnet demonstrated higher coherence and structured responses, while GPT-4o Mini produced concise yet effective outputs. Qwen 2.5 benefited the most from RAG, showing a 14.38% increase in performance, likely due to its stronger reliance on retrieved context for accurate quantum code generation. These results highlight the effectiveness of retrieval-augmented methods in enhancing quantum code generation accuracy and adherence to best practices.
IV Statistical Evaluation of the Dataset
A comprehensive analysis of the dataset was conducted to assess the distribution of instruction lengths, response lengths, response variations across feature groups, and their correlation with quantum functionalities. This section provides insights into the dataset’s structure, response variability, and its implications for training LLMs.
IV-A Response Lengths Across Feature Groups
Figure 5 provides a box plot representation of response lengths across different feature groups.
  • Response lengths remain relatively stable across feature groups, suggesting uniformity in generated outputs.
  • Outliers appear in every feature category, indicating occasional long responses, particularly in core operations, measurement, and mathematical functions.
  • The “Other” category has the shortest responses, likely due to simpler queries requiring minimal output.
Figure 5: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions
IV-B Instruction and Response Length Distributions
Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in the left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost scatter plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 6. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
Figure 6: Feature Group Usage Across Domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories, with color intensity reflecting count.
IV-D Feature Category Association with Quantum Functions
To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 7.
Figure 7: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features. The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns). This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
Quantum gates and operators exhibit the strongest associations with multiple features, particularly with commonly used functions such as qml.PauliZ, qml.RX, and qml.RY, which are integral to quantum circuit construction. Core operations and measurement functions also display high correlations, reflecting their frequent application in quantum circuit execution and result interpretation. Meanwhile, error handling and mathematical functions show relatively lower but still significant interactions with select quantum features, indicating their importance in specific computational scenarios rather than across all use cases.
IV-E Insights and Implications
The findings from Figures 4, 5, 7, and 6 offer several key insights:
  • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
  • Response Length Variability: Most responses are concise, but outliers indicate that certain quantum features demand more detailed explanations.
  • Feature-Specific Trends: Quantum gates and measurement functions dominate dataset interactions, aligning with core PennyLane applications.
  • Domain-Specific Usage: Different quantum fields rely on distinct feature sets, highlighting potential customization opportunities for specialized LLM training.
V Case Studies
The case studies compare base LLM-generated PennyLane code with RAG-enhanced implementations, showcasing notable improvements in functionality, syntax, and modularity across different quantum programming tasks.
In the first task, implementing a 1-qubit parameterized RX gate, the base model introduced redundant imports, suboptimal measurement choices, and missing hardware-specific optimizations. The RAG-optimized implementation addressed these issues by incorporating hardware-aware device initialization, optimized measurement practices, and type annotations, achieving a perfect 10/10 functionality score.
The second task, developing an RX gate circuit with optional parameter binding, revealed static parameter handling and the absence of deferred execution in the base LLM output. The RAG-enhanced version introduced flexible parameter binding, probabilistic measurement strategies, and deferred execution patterns, improving adaptability and modularity.
The third case study, constructing a variational quantum classifier using StronglyEntanglingLayers, further demonstrated the superiority of RAG-based generation. The base LLM solution lacked adaptive training integration, efficient optimization, and TensorFlow-Keras support, requiring manual gradient updates. In contrast, the RAG-optimized model seamlessly integrated TensorFlow, Keras, and the Adam optimizer, significantly enhancing training efficiency and parameter optimization.
Overall, across all tasks, RAG-assisted code generation consistently outperformed the base LLM, producing more structured, modular, and functionally superior PennyLane code. This highlights the effectiveness of retrieval-augmented approaches in enhancing quantum programming workflows.
Task: Implement 1-qubit parameterized RX gate using PennyLane
Base LLM Implementation (Without RAG) Limitations Identified: Unnecessary numpy import Suboptimal measurement choice (expectation value vs state) Missing hardware-specific optimizations 1import pennylane as qml 2from pennylane import numpy as np # Redundant import 3 4dev = qml.device(’default.qubit’, wires=1) 5 6@qml.qnode(dev) 7def circuit(theta): 8 qml.RX(theta, wires=0) 9 return qml.expval(qml.PauliZ(0)) # Non-optimal measurement Listing 1: Non-RAG Implementation Evaluation Results: Functionality: 9/10 (Valid but limited measurement) Syntax: 9/10 (Redundant imports) Modularity: 8/10 (Missing type hints)
RAG-Enhanced Implementation Retrieved Context: Hardware-aware device initialization (sample_1137.txt) State measurement best practices (sample_214.txt) Testing methodologies (sample_23.txt) 1import pennylane as qml 2 3dev = qml.device("lightning.qubit", wires=1) # Hardware-optimized 4 5@qml.qnode(dev) 6def circuit(theta: float): # Type hinting 7 qml.RX(theta, wires=0) 8 return qml.state() # Full state measurement Listing 2: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Complete state analysis) Syntax: 10/10 (Clean imports) Modularity: 9/10 (Type annotations)
Task: Create 1-qubit RX gate circuit with optional parameter binding
Base LLM Implementation (Without RAG) Limitations Identified: Static parameter handling Missing deferred parameterization Limited measurement flexibility 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 if theta: # Binary parameter check 6 qml.RX(theta, wires=0) 7 return qml.expval(qml.PauliZ(0)) # Fixed measurement Listing 3: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (No variable binding) Syntax: 9/10 (Valid but rigid) Modularity: 8/10 (No late binding support)
RAG-Enhanced Implementation Retrieved Context: Variable parameter techniques (sample_1137.txt) Probabilistic measurement strategies (sample_214.txt) Deferred execution patterns (sample_23.txt) 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 param = theta if theta else qml.variable(’theta’) 6 qml.RX(param, wires=0) 7 return qml.probs(wires=0) # Full probability measurement Listing 4: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Flexible parameter binding) Syntax: 9/10 (Clear variable handling) Modularity: 9/10 (Supports late binding)
Task: Implement a Variational Quantum Classifier with Strongly Entangling Layers
Base LLM Implementation (Without RAG) Limitations Identified: Uses a static quantum circuit with fixed parameters, lacking adaptive training integration. Lacks integration with TensorFlow/Keras, requiring a manual optimization loop. Uses Gradient Descent Optimizer instead of the Adam optimizer, which may slow convergence. 1import pennylane as qml 2import numpy as np 3from sklearn.datasets import make_classification 4from sklearn.model_selection import train_test_split 5from sklearn.preprocessing import StandardScaler 6from sklearn.metrics import accuracy_score 7 8# Generate a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Scale the features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define the quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Define the quantum circuit 21@qml.qnode(dev) 22def circuit(params, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.StronglyEntanglingLayers(params, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 5: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (Works but lacks adaptive training integration) Syntax: 9/10 (Code is valid but manual optimization is not ideal) Modularity: 8/10 (Training and inference should be better structured)
RAG-Enhanced Implementation Retrieved Context: TensorFlow-Keras integration for efficient optimization. Adam optimizer instead of manual updates for improved convergence. Improved quantum circuit design based on best practices. 1import pennylane as qml 2import numpy as np 3import tensorflow as tf 4from sklearn.datasets import make_classification 5from sklearn.model_selection import train_test_split 6from sklearn.preprocessing import StandardScaler 7 8# Create a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Standardize features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Quantum circuit with Strongly Entangling Layers 21@qml.qnode(dev) 22def quantum_circuit(weights, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.templates.StronglyEntanglingLayers(weights, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 6: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Seamless Keras integration for training) Syntax: 10/10 (Concise and efficient TensorFlow pipeline) Modularity: 9/10 (Well-structured training process)
VI Conclusion
This paper presents a dedicated PennyLane dataset and a systematic methodology for preparing quantum code snippets for LLM training. Our open-source release aims to empower researchers and developers to build advanced quantum code assistants that reduce onboarding difficulties and accelerate development in hybrid quantum-classical systems. Ongoing work includes expanding the dataset with more domain-specific samples, conducting broader evaluations with different LLMs, and exploring cross-framework synergy for universal quantum code generation.
Acknowledgment
This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and Center for CyberSecurity (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104.

### v2

@@
 Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning.
-Their potential in aiding quantum software development remains underexplored, particularly for the PennyLane framework—a leading platform for hybrid quantum-classical computing.
-To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific code samples of quantum circuits and their contextual descriptions, specifically curated to train/fine-tune LLM-based quantum code assistance.
-Our key contributions are threefold: (1) the automatic creation and open-source release of a comprehensive PennyLane dataset leveraging quantum computing textbooks, official documentation, and open-source repositories; (2) the development of a systematic methodology for data refinement, annotation, and formatting to optimize LLM training efficiency; and (3) a thorough evaluation, based on a Retrieval-Augmented Generation (RAG) framework, demonstrating the effectiveness of our dataset in streamlining PennyLane code generation and improving quantum development workflows.
-Compared to existing efforts that predominantly focus on Qiskit, our dataset significantly broadens the spectrum of quantum frameworks covered in AI-driven code assistance.
-By bridging this gap and providing reproducible dataset-creation methodologies, we aim to advance the field of AI-assisted quantum programming, making quantum computing more accessible to both newcomers and experienced developers.
+However, their application in quantum software development remains underexplored, particularly for PennyLane—a leading framework for hybrid quantum-classical computing.
+To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific quantum code samples and contextual descriptions, specifically curated to support LLM training and fine-tuning for quantum code assistance.
+Our contributions are threefold: (1) the automatic construction and open-source release of a comprehensive PennyLane dataset derived from textbooks, official documentation, and open-source repositories; (2) a structured methodology for data curation, annotation, and formatting to enhance LLM usability and relevance; and (3) a rigorous evaluation of code generation capabilities using both baseline Retrieval-Augmented Generation (RAG) and a GraphRAG-enhanced pipeline.
+Using the PennyLang framework, we demonstrate that GraphRAG, when applied to a GPT-4o Mini model, substantially outperforms standard prompting and baseline RAG.
+Accuracy improves from 20.5% (without RAG) to 58.2% with GraphRAG, showcasing its effectiveness in reducing hallucinations and improving code correctness in quantum programming tasks.
+Compared to prior efforts focused largely on Qiskit, our work expands LLM-based assistance to the PennyLane ecosystem, contributing practical tools and reproducible methodologies for advancing AI-assisted quantum software development.
 Index Terms:
-Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), Code Completion and Assistance, Dataset Curation and Annotation, Evaluation Framework
+Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
 I Introduction
-Quantum computing has emerged as a transformative technology, with the potential to solve complex problems that are infeasible for classical systems [1, 2].
-Breakthroughs such as Google’s demonstration of “quantum supremacy” [3]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [4, 5, 6, 7, 8].
-Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks, hindering widespread adoption [9, 10].
+Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3].
+Breakthroughs such as Google’s demonstration of “quantum supremacy” [4]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
+Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [16, 17, 18].
 PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19].
-However, unlike the Qiskit framework [12], which benefits from dedicated AI-driven code assistants [13], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
-In parallel, the field of Natural Language Processing (NLP) has witnessed a surge in the capabilities of LLMs, which excel at reasoning, contextual understanding, and automated code generation [14, 15, 16].
-These models have been successfully leveraged in classical programming [17, 18], improving developer efficiency and reducing syntactical and logical errors.
-However, their application in quantum programming remains largely underexplored, particularly for PennyLane.
-This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for its refinement, annotation, formatting, and evaluation to facilitate LLM training.
-State-of-the-Art and Limitations: While several studies have explored LLM-driven code generation for classical programming tasks [19, 20], research on quantum-specific datasets remains scarce.
-The most notable effort in this domain is the Qiskit Code Assistant [13], which leverages a Qiskit-specific dataset for AI-assisted quantum code generation.
-However, its hardware-centric approach does not extend to PennyLane, which specializes in variational quantum circuits and quantum machine learning applications [21].
+However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
+In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25].
+These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors.
+However, their application in quantum programming—especially within the PennyLane ecosystem—remains underexplored.
+This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
+State-of-the-Art and Limitations: While LLMs have been benchmarked extensively for classical programming [28, 29], their performance in quantum domains is less understood.
+Qiskit HumanEval and Qiskit Code Assistant [22] represent early attempts to extend LLM capabilities into quantum programming, but they focus primarily on Qiskit—a hardware-centric framework.
+PennyLane, in contrast, is designed for variational quantum algorithms and quantum machine learning [30], requiring a different kind of LLM support and training corpus.
 The primary challenge in developing a PennyLane code assistant lies in data availability.
@@
 Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
-Our Approach and Contributions: In this work, we introduce PennyLang to address the gap in LLM-assisted quantum code generation for pennylane through the following novel contributions.
-An overview of our contributions is presented in Fig.
-1.
+Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset tailored for PennyLane code generation, along with a rigorous evaluation pipeline.
+In addition to a baseline RAG framework, we propose and test the GraphRAG framework [32], which leverages structured code relationships to improve the quality and relevance of retrieval during generation.
+Our contributions are summarized in Fig.
+1 and detailed below:
 Figure 1: Overview of novel contributions in this work.
   1. 1.
-    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, two quantum computing textbooks, and the official PennyLane documentation.
-Each sample is meticulously annotated with explanatory comments and contextual descriptions, facilitating more effective LLM training for quantum programming.
+    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, quantum computing textbooks, and official PennyLane documentation.
+Each sample is enriched with contextual descriptions and annotations to support instruction tuning.
   2. 2.
-    Systematic Data Refinement and Annotation: We develop a robust methodology for dataset cleaning, eliminating non-PennyLane snippets and enriching the samples with structured metadata.
-To further enhance usability, we employ the GPT-4o API to convert code samples into an instruction-query format, providing clear guidance for model training in a structured, task-oriented manner.
+    Systematic Data Refinement and Annotation: We present a robust methodology for dataset cleaning and metadata enrichment.
+Using the GPT-4o API, code snippets are transformed into instruction-query pairs, improving compatibility with LLM training pipelines.
   3. 3.
@@
 The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
-Summary of Key Results: RAG-enhanced quantum code generation significantly improved performance across all models.
+  4. 4.
+    GraphRAG-Enhanced Evaluation Framework: We extend the baseline RAG framework with GraphRAG, which introduces a graph-based context retrieval mechanism for code augmentation.
+GraphRAG models contextual and structural relationships between quantum code samples, enabling more accurate and consistent retrievals.
+We evaluate this setup using GPT-4o Mini model, comparing GraphRAG, and no-RAG scenarios in a controlled setting.
+Summary of Key Results: RAG-enhanced quantum code generation improves performance across all models.
 GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context.
 These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability.
+Our results from GraphRAG significantly outperforms both baseline RAG and vanilla prompting.
+When applied to GPT-4o Mini, GraphRAG boosts accuracy from 20.5% (no RAG) to 58.2%, marking a 37.7% absolute improvement in correct solutions across PennyLang dataset, confirming the benefit of structured retrieval in quantum programming tasks.
 Paper Organization: Sec.
-II provides background information and reviews related work.
+II reviews related work and background on quantum programming and RAG.
 Sec.
-III outlines our methodology for dataset collection, cleaning, and annotation along with the tokenization and formatting pipeline used for preprocessing.
+III outlines our dataset construction, annotation methodology, and the GraphRAG retrieval pipeline.
 Sec.
-IV presents experimental results, including an evaluation of our dataset using a RAG-based framework.
+IV presents experimental results comparing model performance across prompting settings.
 Sec.
-V presents some case studies Finally, Sec.
-VI concludes the paper with a summary of findings and future research directions.
+V discusses case studies and qualitative insights.
+Sec.
+VI concludes with future directions for dataset-driven quantum code assistance.
 II Background and Related Work
-TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
+TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training.
 Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [33] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [34] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [35] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [36] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [37] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [38] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [39] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [40] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
 The intersection of LLMs and quantum computing is an emerging research domain.
-While LLMs have demonstrated impressive capabilities in NLP [31] and automated code generation [32], their application in quantum software development remains under-explored.
-Existing studies on AI-driven code assistants primarily focus on classical programming languages such as Python, Java, and C++, or quantum computing frameworks like Qiskit [13].
-However, the PennyLane framework [11], widely used for hybrid quantum-classical computing, lacks dedicated research efforts in AI-assisted quantum programming.
+While LLMs have demonstrated impressive capabilities in NLP [41] and automated code generation [42], their application in quantum software development remains underexplored.
+Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows.
 This section presents an in-depth review of related work in three key areas:
@@
   • CodeLlama-Quantum (Meta AI) [27] extended CodeLlama for Qiskit-based applications.
-Despite these advancements, challenges persist:
+However, these works face key limitations:
   1. 1.
-    Limited Training Data: Quantum programming lacks large-scale, high-quality datasets.
+    Lack of High-Quality Data: Quantum programming lacks scalable, curated training datasets.
   2. 2.
-    Complexity of Quantum Circuits: Unlike classical code, quantum programs require understanding quantum mechanics.
+    Code Complexity: Quantum circuits require LLMs to reason over non-trivial entanglement and measurement logic.
   3. 3.
-    Hybrid Workflows: PennyLane integrates classical and quantum processing, requiring LLMs to learn hybrid paradigms.
-These challenges underscore the need for PennyLane-specific datasets for LLM training, as none currently exist, reinforcing the importance of our research.
+    Hybrid Workflows: PennyLane introduces mixed classical-quantum code patterns that require deeper contextual understanding.
+To bridge this gap, we propose the first PennyLane-centric dataset—PennyLang—paired with an advanced evaluation pipeline that incorporates RAG-based retrieval.
+II-C Retrieval-Augmented Generation and GraphRAG
+Retrieval-Augmented Generation (RAG) [47] has emerged as a popular method to ground LLM outputs using external context.
+In quantum programming, RAG helps models compensate for limited training data and outdated knowledge by retrieving relevant examples at inference time.
+However, baseline RAG systems suffer from semantic drift, irrelevant context injection, and weak code relevance due to shallow vector-based similarity matching.
+To address this, we employ GraphRAG[32], a graph-enhanced retrieval framework designed to capture structural and functional relationships between quantum code samples.
+GraphRAG models PennyLang as a graph of interconnected examples, allowing retrievals that are topologically and semantically aligned with the user query.
+Compared to traditional RAG, GraphRAG significantly improves context precision and reduces hallucinations in generated code.
+This leads to large gains in generation accuracy on the PennyLang test dataset.
 Figure 2: Methodology for PennyLang quantum code generation and evaluation.
@@
 TABLE II: Comparison of Related Work and Our Contributions
-Research Area Existing Works Our Contributions
+Research Area Existing Works Our Contributions.
 LLMs for Code Generation Codex [26], CodeBERT [43] PennyLane-specific dataset
@@
   • Quantum chemistry applications.
-TABLE III: Tokenized Input and Attention Masking
+TABLE III: Tokenized Input and Attention Masking.
 Tokenized Input Attention Mask
@@
 The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
-TABLE IV: Dataset Composition
+TABLE IV: Dataset Composition.
 Source Samples Collected
@@
 This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
-TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
+TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework.
 Component Function Description
@@
 ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
-TABLE VI: LLMs Used in the RAG-Based Evaluation Framework
+TABLE VI: LLMs Used in the RAG-Based Evaluation Framework.
 Model Parameters Architecture Provider Usage in Pipeline
@@
   • Impact on Training: The self-attention mechanism [52] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
+Figure 3: Distribution of instruction and response lengths with their correlation.
+The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
 III-C Methodology for Dataset Evaluation
 III-C1 Retrieval-Augmented Generation (RAG) Framework
-To enhance LLM-based quantum code generation, we employ a Retrieval-Augmented Generation (RAG) methodology that integrates retrieval-based context with language model inference.
-This hybrid approach significantly improves:
-  • Accuracy: By grounding responses in retrieved dataset samples, reducing model hallucination.
-  • Adherence to PennyLane’s best practices: Ensuring generated quantum code aligns with official standards.
-  • Response Quality: Structured retrieval enriches contextual understanding, improving coherence and correctness.
+To assess the impact of context-aware prompting on LLM-based quantum code generation, we adopt a Retrieval-Augmented Generation (RAG) framework.
+This method enhances generation quality by incorporating relevant external context retrieved from our curated PennyLang dataset.
+RAG helps address common shortcomings in LLMs, such as hallucinations and outdated or incorrect syntax, by anchoring their output in concrete examples from real-world PennyLane programs.
+This results in improved syntactic validity, functional correctness, and alignment with PennyLane best practices.
 III-C2 RAG Implementation using LangChain
-LangChain is utilized to construct a modular and efficient RAG pipeline. The framework consists of the following core components:
-  1. 1.
-    Dataset Loading: PennyLane dataset samples are extracted from curated dataset structured in intruction-response format.
-  2. 2.
-    Embedding Generation: Each sample is tokenized and converted into high-dimensional vector embeddings using OpenAIEmbeddings.
-  3. 3.
-    Vector Storage and Retrieval: The embeddings are stored in a Chroma vector database for fast retrieval.
-  4. 4.
-    Prompting and Response Generation:
-    • When a query is submitted, Maximum Marginal Relevance (MMR) search fetches the most relevant samples in RAG-augmented input.
-    • Retrieved context is formatted into a structured prompt.
-    • The LLM model generates responses using both RAG-augmented and non-RAG prompts.
-  5. 5.
-    Evaluation: The responses are compared using a second LLM (GPT-4o Mini), which scores them based on:
-    • Functionality: Logical correctness of the quantum circuit.
-    • Syntax: Validity of the PennyLane code.
-    • Modularity: Adherence to best practices in structuring quantum workflows.
-Figure 3: Performance of GPT 4o mini, Claude, and Qwen With and Without RAG.
-Mean scores (±1 SD) over 29 trials for GPT 4o mini, Claude, and Qwen under both conditions (blue: With RAG; orange: Without RAG), with individual scores shown as swarmplot markers.
-Figure 3 presents the results of our RAG-based evaluation methodology, demonstrating a significant performance improvement with retrieval-augmented responses.
-Each model was tested with 30 code samples, and responses were evaluated using GPT-4o, which graded them based on syntax, functionality, and modularity.
-The results indicate that the RAG-enhanced approach consistently outperforms the base LLM, yielding higher overall scores.
-III-C3 LLM Models and Configuration
-The models used in this framework differ in size, architecture, and inference capabilities.
-Table VI provides a comprehensive comparison of key characteristics including parameter count, context window size, and quantum-specific training data percentage.
-Figure 4: Distribution of instruction and response lengths with their correlation.
-The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
-III-C4 Prompt Design for RAG and Non-RAG Evaluation
-To ensure fair evaluation, we construct two different prompts—one for RAG-augmented inference and another for direct model inference without retrieval.
-  • Non-RAG Prompt:
-    Base LLM Prompt (Without RAG) "You are an advanced assistant for quantum computing tasks using PennyLane.
-Provide a response to the user’s question.
-Make sure to follow best practices and optimize for efficiency.
-If you are unsure, state that clearly."
-  • RAG-Enhanced Prompt:
-    RAG-Enhanced Prompt with Context Retrieval "You are an advanced assistant for quantum computing tasks using PennyLane.
-Use the following pieces of retrieved context to answer the question.
-Use multiple retrieved code snippets whenever necessary to improve accuracy.
-If a single snippet does not provide enough information, retrieve and combine relevant parts from multiple sources.
-Make sure to follow best practices and optimize for efficiency.
-If you don’t know the answer, say that you don’t know." "{context}"
-III-C5 Evaluation Criteria and Scoring
-To assess the effectiveness of RAG-based retrieval, generated responses are evaluated using GPT-4o with the following scoring prompt:
-Evaluation Prompt for GPT-4o You are an AI assistant tasked with evaluating two different answers to the same question.
-Question: {question} Answer with RAG: {rag_answer} Answer without RAG: {non_rag_answer} Please provide a score from 1 to 30 for each answer based on: - Functionality (correctness of quantum operations) - Syntax (validity of PennyLane code) - Modularity (code organization and efficiency) Justify your scoring.
-III-C6 Findings and Impact of RAG Integration
-After evaluating multiple quantum programming queries using the RAG vs.
-Non-RAG framework, we observe: Based on the evaluation results, RAG-enhanced responses consistently outperformed non-RAG responses across all tested models—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5.
-On average, RAG integration improved response scores by 11.67% in GPT-4o Mini, 7.69% in Claude 3.5, and 14.38% in Qwen 2.5.
-The improvement was most noticeable in functionality and syntax, as retrieved context helped prevent incorrect quantum operations and improved PennyLane-specific syntax accuracy.
-Claude 3.5 Sonnet demonstrated higher coherence and structured responses, while GPT-4o Mini produced concise yet effective outputs.
-Qwen 2.5 benefited the most from RAG, showing a 14.38% increase in performance, likely due to its stronger reliance on retrieved context for accurate quantum code generation.
-These results highlight the effectiveness of retrieval-augmented methods in enhancing quantum code generation accuracy and adherence to best practices.
+Our RAG pipeline is implemented using the LangChain framework, which provides a modular infrastructure for document loading, vector embedding, retrieval, and prompt formatting.
+The dataset—formatted in an instruction-response structure—is first embedded using the OpenAI Embeddings module.
+The resulting vectors are stored in a Chroma database for efficient retrieval.
+When a user query is submitted (e.g., a quantum programming task or challenge description), we apply Maximum Marginal Relevance (MMR) search to extract the top relevant examples from the dataset.
+These examples are concatenated into a structured prompt and fed into the LLM for code generation.
+To ensure fairness and evaluate RAG’s effectiveness, we generate responses both with and without retrieved context using the same query.
+III-C3 RAG Evaluation
+To benchmark the performance of RAG-augmented LLMs, we evaluate three state-of-the-art models: GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5-7B-Instruct-Turbo.
+Each model is tested on 29 quantum code generation tasks sampled from the QHack challenge dataset.
+For each query, both a baseline (non-RAG) and a RAG-enhanced response is generated.
+We then use GPT-4o to evaluate the outputs according to a rubric comprising three criteria:
+  • Functionality: Does the code produce the correct quantum behavior?
+  • Syntax: Is the code valid and executable in PennyLane?
+  • Modularity: Is the solution well-organized and reusable?
+Figure 4: Performance of GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5 with and without RAG. Mean scores (±1 SD) over 29 trials are shown for each model. Blue: RAG; Orange: Non-RAG.
+Figure 4 summarizes the outcome of these evaluations.
+RAG-enhanced prompting consistently improves model performance across all criteria.
+On average, GPT-4o Mini improves by 11.67%, Claude 3.5 by 7.69%, and Qwen 2.5 by 14.38%.
+These gains highlight how retrieval provides critical context, especially for complex or multi-step quantum circuits.
+III-C4 GraphRAG for Advanced Retrieval
+To further improve retrieval precision beyond vector-based similarity, we extend our RAG framework using GraphRAG.
+While the baseline RAG implementation relies solely on Maximum Marginal Relevance (MMR) vector retrieval, GraphRAG incorporates structural relationships between code samples to guide context selection.
+In this method, the PennyLang dataset is organized into a graph where nodes represent code samples and edges represent functional or structural similarity—such as shared quantum gates, observables, and template usage patterns.
+When a query is received, GraphRAG identifies a subgraph around the most semantically similar samples and performs hybrid retrieval by combining graph neighborhood traversal with traditional vector similarity.
+This enables the selection of more functionally aligned and topologically relevant code snippets, resulting in higher quality context for LLMs.
+We evaluate GraphRAG using the same task set as our baseline experiments but restrict the evaluation to the GPT-4o Mini model due to its superior baseline performance.
+The model was tested on 146 PennyLang test tasks in both no-RAG and GraphRAG settings.
+The correctness of the generated outputs was determined by whether the solution passed the associated unit tests.
+TABLE VII: Performance of GPT-4o Mini: No-RAG vs. GraphRAG (146 tasks).
+Metric No-RAG GraphRAG
+Correct Solutions 30 85
+Incorrect Solutions 116 61
+Total Samples 146 146
+Accuracy (%) 20.55% 58.22%
+As shown in Table VII, the GraphRAG approach yields a dramatic improvement in accuracy—from 20.55% without retrieval to 58.22% with graph-enhanced retrieval.
+This threefold increase highlights the value of graph-based structural reasoning in quantum programming tasks, particularly for problems involving complex circuit construction or uncommon PennyLane APIs.
+Together, these results establish GraphRAG as a significant enhancement to conventional RAG pipelines and demonstrate its practical utility in improving the performance of LLMs on challenging real-world quantum programming benchmarks.
 IV Statistical Evaluation of the Dataset
@@
 Figure 5: Box plot of response lengths across different feature groups.
-Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions
+Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions.
 IV-B Instruction and Response Length Distributions
@@
 VI Conclusion
-This paper presents a dedicated PennyLane dataset and a systematic methodology for preparing quantum code snippets for LLM training.
-Our open-source release aims to empower researchers and developers to build advanced quantum code assistants that reduce onboarding difficulties and accelerate development in hybrid quantum-classical systems.
-Ongoing work includes expanding the dataset with more domain-specific samples, conducting broader evaluations with different LLMs, and exploring cross-framework synergy for universal quantum code generation.
+In this paper, we introduce PennyLang, a high-quality, instruction-response dataset tailored for PennyLane-based quantum code generation.
+We propose a robust data collection and annotation methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples.
+To assess the practical utility of the dataset, we develop a Retrieval-Augmented Generation (RAG) evaluation framework and benchmark three state-of-the-art LLMs—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B—under both RAG and non-RAG conditions.
+Our findings demonstrate consistent performance gains with RAG, especially in functionality and syntax correctness.
+Building on this, we introduce a GraphRAG extension that models structural relationships across code snippets using graph-based retrieval.
+When applied to GPT-4o Mini, GraphRAG yields a 2.8$\times$ improvement in accuracy compared to the non-RAG baseline, highlighting its potential for deeper semantic retrieval in quantum programming.
+We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development.
+Future directions include training lightweight instruction-tuned models on this dataset, scaling GraphRAG across different models and quantum frameworks, and expanding toward cross-framework generalization between PennyLane, Qiskit, and Cirq.
 Acknowledgment
 This work was supported in part by the NYUAD Center for Quantum and Topological Systems (CQTS), funded by Tamkeen under the NYUAD Research Institute grant CG008, and Center for CyberSecurity (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104.
+This research was carried out on the High Performance Computing resources at New York University Abu Dhabi.

### v3

@@
 Abstract
-Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning.
-However, their application in quantum software development remains underexplored, particularly for PennyLane—a leading framework for hybrid quantum-classical computing.
-To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific quantum code samples and contextual descriptions, specifically curated to support LLM training and fine-tuning for quantum code assistance.
-Our contributions are threefold: (1) the automatic construction and open-source release of a comprehensive PennyLane dataset derived from textbooks, official documentation, and open-source repositories; (2) a structured methodology for data curation, annotation, and formatting to enhance LLM usability and relevance; and (3) a rigorous evaluation of code generation capabilities using both baseline Retrieval-Augmented Generation (RAG) and a GraphRAG-enhanced pipeline.
-Using the PennyLang framework, we demonstrate that GraphRAG, when applied to a GPT-4o Mini model, substantially outperforms standard prompting and baseline RAG.
-Accuracy improves from 20.5% (without RAG) to 58.2% with GraphRAG, showcasing its effectiveness in reducing hallucinations and improving code correctness in quantum programming tasks.
-Compared to prior efforts focused largely on Qiskit, our work expands LLM-based assistance to the PennyLane ecosystem, contributing practical tools and reproducible methodologies for advancing AI-assisted quantum software development.
+Large Language Models (LLMs) offer powerful capabilities in code generation, natural language understanding, and domain-specific reasoning.
+Their application to quantum software development remains limited, in part because of the lack of high-quality datasets both for LLM training and as dependable knowledge sources.
+To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories.
+Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline.
+Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness.
+Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
 Index Terms:
@@
 Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3].
-Breakthroughs such as Google’s demonstration of “quantum supremacy” [4]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
-Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [16, 17, 18].
+Breakthroughs such as Google’s demonstration of “quantum supremacy” [4], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
+Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [17, 18].
 PennyLane is a leading open-source Python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19].
-However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
+However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers in writing optimized quantum code.
 In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25].
 These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors.
-However, their application in quantum programming—especially within the PennyLane ecosystem—remains underexplored.
+However, their application in quantum programming, especially within the PennyLane ecosystem, remains underexplored.
 This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
-State-of-the-Art and Limitations: While LLMs have been benchmarked extensively for classical programming [28, 29], their performance in quantum domains is less understood.
-Qiskit HumanEval and Qiskit Code Assistant [22] represent early attempts to extend LLM capabilities into quantum programming, but they focus primarily on Qiskit—a hardware-centric framework.
-PennyLane, in contrast, is designed for variational quantum algorithms and quantum machine learning [30], requiring a different kind of LLM support and training corpus.
-The primary challenge in developing a PennyLane code assistant lies in data availability.
-Unlike classical programming languages, where vast repositories of well-documented code exist, quantum computing resources are fragmented across research papers, user forums, open-source repositories, and official documentation [31].
-Moreover, raw quantum code samples often lack contextual annotations, making it difficult for LLMs to learn meaningful associations between quantum operations and their intended applications.
-Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset tailored for PennyLane code generation, along with a rigorous evaluation pipeline.
-In addition to a baseline RAG framework, we propose and test the GraphRAG framework [32], which leverages structured code relationships to improve the quality and relevance of retrieval during generation.
-Our contributions are summarized in Fig.
-1 and detailed below:
+State-of-the-Art and Limitations: LLMs have been extensively benchmarked for classical programming [28, 29], but their performance in quantum domains remains less explored.
+Early efforts such as Qiskit HumanEval and Qiskit Code Assistant [22] adapt LLMs to quantum programming but focus on Qiskit’s hardware-centric paradigm.
+PennyLane, by contrast, targets variational quantum algorithms and quantum machine learning [30], requiring dedicated LLM support and a specialized training corpus.
+A major barrier is the scarcity of suitable data: quantum code is scattered across papers, forums, repositories, and documentation [31], and often lacks the contextual annotations necessary for LLMs to link quantum operations with their intended use.
 Figure 1: Overview of novel contributions in this work.
+Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset specifically designed for PennyLane code generation, accompanied by a rigorous evaluation pipeline.
+We implement a baseline Retrieval-Augmented Generation (RAG) framework to assess the dataset’s effectiveness across multiple models.
+Our key contributions, shown in Fig.
+1, consist of:
   1. 1.
-    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, quantum computing textbooks, and official PennyLane documentation.
-Each sample is enriched with contextual descriptions and annotations to support instruction tuning.
+    Comprehensive PennyLane Dataset: Curated 3,347 PennyLane-specific code samples from GitHub repositories, quantum computing textbooks, and official PennyLane documentation, each enriched with contextual descriptions and annotations to support instruction tuning.
   2. 2.
-    Systematic Data Refinement and Annotation: We present a robust methodology for dataset cleaning and metadata enrichment.
-Using the GPT-4o API, code snippets are transformed into instruction-query pairs, improving compatibility with LLM training pipelines.
+    Systematic Data Refinement and Annotation: Developed a robust pipeline for dataset cleaning, metadata enrichment, and conversion of code snippets into instruction–query pairs using the GPT-4o API, ensuring compatibility with LLM training workflows.
   3. 3.
-    RAG-Based Evaluation Framework: We use a Retrieval-Augmented Generation (RAG) evaluation framework to assess the dataset’s impact on model performance.
-The framework compares outputs from three models—GPT-4o mini, Claude 3.5, and Qwen 7B—under two conditions: (i) with access to the dataset as retrieval-augmented context and (ii) without dataset augmentation.
-The generated outputs are validated and scored based on functionality, syntax correctness, and modularity.
-  4. 4.
-    GraphRAG-Enhanced Evaluation Framework: We extend the baseline RAG framework with GraphRAG, which introduces a graph-based context retrieval mechanism for code augmentation.
-GraphRAG models contextual and structural relationships between quantum code samples, enabling more accurate and consistent retrievals.
-We evaluate this setup using GPT-4o Mini model, comparing GraphRAG, and no-RAG scenarios in a controlled setting.
-Summary of Key Results: RAG-enhanced quantum code generation improves performance across all models.
-GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context.
-These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability.
-Our results from GraphRAG significantly outperforms both baseline RAG and vanilla prompting.
-When applied to GPT-4o Mini, GraphRAG boosts accuracy from 20.5% (no RAG) to 58.2%, marking a 37.7% absolute improvement in correct solutions across PennyLang dataset, confirming the benefit of structured retrieval in quantum programming tasks.
-Paper Organization: Sec.
-II reviews related work and background on quantum programming and RAG.
-Sec.
-III outlines our dataset construction, annotation methodology, and the GraphRAG retrieval pipeline.
-Sec.
-IV presents experimental results comparing model performance across prompting settings.
-Sec.
-V discusses case studies and qualitative insights.
-Sec.
-VI concludes with future directions for dataset-driven quantum code assistance.
+    RAG-Based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance.
+Evaluated GPT-4o mini, Claude 3.5, Qwen 7B, and LLaMa 4 Maverick 17B under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
+Summary of Key Results: Retrieval augmentation with PennyLang yields substantial relative gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66% ($\approx$32.95% relative increase), and LLaMa 4 improves from 78.78% to 84.84% ($\approx$7.7% relative increase).
+In contrast, stronger baselines (GPT-4o-mini and Claude 3.5 Sonnet) do not benefit from naive full-context retrieval, their best performance is achieved without retrieval (85.60% and 95.07%, respectively), with full retrieval resulting in slight degradation.
+Detailed analysis is presented in Section 5.
 II Background and Related Work
-TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training.
-Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Textbook Dataset [33] Qiskit Qiskit.org Textbook Multiple structured examples Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [34] Qiskit IBM Quantum Experience Derived from past challenges Instruction-Response Format JSON/Notebook Qiskit Code Repository [35] Qiskit GitHub (Qiskit AI repos) Thousands (Scattered) Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [36] Qiskit Kaggle (Possible) 309 (Limited samples) Pretrained Benchmark Data CSV/JSON QDataSet [37] Qiskit GitHub 52 (raw datasets) Quantum ML Data Various QCircuitNet [38] Qiskit Research Paper 4,800 Samples Quantum Algorithm Design Benchmark Python/QASM PennyLane Documentation [39] PennyLane PennyLane.ai Docs Various examples Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [40] PennyLane GitHub Open-Source Scattered code samples Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Instruction-Response Dataset Curated and Structured Dataset JSON file
+TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
+Dataset Name Framework Source Samples Available Dataset Type Format Qiskit Framework Datasets Qiskit Textbook Dataset [32] Qiskit Qiskit.org Textbook Multiple Structured Qiskit Code Python Scripts IBM Quantum Challenge Data [33] Qiskit IBM Quantum Experience Derived Instruction-Response Format JSON/Notebook Qiskit Code Repository [34] Qiskit GitHub (Qiskit AI repos) Thousands Raw Qiskit Code Samples Python Scripts Q-Gen Quantum Circuit [35] Qiskit Kaggle (Possible) 309 Pretrained Benchmark Data CSV/JSON QDataSet [36] Qiskit GitHub 52 Quantum ML Data Various QCircuitNet [37] Qiskit Research Paper 4,800 Quantum Algorithm Design Benchmark Python/QASM PennyLane Framework Datasets PennyLane Documentation [38] PennyLane PennyLane.ai Docs Various Raw PennyLane Code Samples Python Scripts PennyLane GitHub Repos [39] PennyLane GitHub Open-Source Scattered Unstructured Quantum Code Python Scripts PennyLang (Ours) PennyLane Open-Source Dataset 3,347 Curated and Structured Dataset JSON file
 The intersection of LLMs and quantum computing is an emerging research domain.
@@
 II-A LLMs for Code Generation in Classical Programming
-Transformer-based models have significantly impacted AI-driven programming, automating code generation, completion, and debugging. Key contributions include:
-  • Codex (OpenAI) [26] introduced a language model trained on public code repositories, enabling natural language-to-code translation.
-  • CodeBERT [43] leveraged bidirectional transformers to enhance source code understanding.
-  • InCoder [29] developed a multi-turn interactive model for code infilling and synthesis.
-  • AlphaCode [25] showed competitive programming-level LLM capabilities.
-Additionally, benchmarking studies have evaluated LLM performance on programming tasks:
-  • HumanEval [24] proposed an execution-based evaluation for Python code generation.
-  • MBPP [44] introduced a multi-language benchmark covering correctness and execution constraints.
-  • CodeXGLUE [45] aggregated multiple datasets to benchmark code-related tasks.
-These studies highlight the necessity of well-structured datasets to enhance LLM code generation performance.
+Transformer-based language models have revolutionized automated code generation, significantly advancing classical software development.
+OpenAI’s Codex [26], trained on large-scale public code repositories, enabled natural language-to-code translation, setting a new standard for programming assistants.
+CodeBERT [42] further improved source code understanding through bidirectional transformers, while InCoder [29] introduced multi-turn interactive capabilities for code infilling and synthesis.
+AlphaCode [25] demonstrated that LLMs could generate solutions for competitive programming problems, reaching near-human performance.
+To rigorously evaluate these models, benchmarks such as HumanEval [24], MBPP [43], and CodeXGLUE [44] have been proposed, each offering complementary perspectives on correctness, diversity, and execution fidelity.
+These benchmarks collectively emphasize the importance of well-curated, structured datasets in improving LLM-driven code generation.
+TABLE II: Comparison: Related Work & Our Contributions
+Research Area Existing Works Our Contributions LLMs for Code Generation Codex [26] CodeBERT [42] PennyLane-specific dataset LLMs for Quantum Programming Qiskit Assistant [22] [45] Hybrid quantum- classical dataset Quantum Code Datasets QCircuitNet [37] QDataset [37] First PennyLane dataset
 II-B LLMs for Quantum Programming
-Existing research on AI-powered quantum code assistants focuses primarily on the Qiskit framework:
-  • Qiskit Code Assistant (IBM Research) [22] assists users in debugging and optimizing Qiskit code.
-  • Qiuchi et al. [46] adapted BERT-based models to improve quantum code understanding.
-  • CodeLlama-Quantum (Meta AI) [27] extended CodeLlama for Qiskit-based applications.
-However, these works face key limitations:
-  1. 1.
-    Lack of High-Quality Data: Quantum programming lacks scalable, curated training datasets.
-  2. 2.
-    Code Complexity: Quantum circuits require LLMs to reason over non-trivial entanglement and measurement logic.
-  3. 3.
-    Hybrid Workflows: PennyLane introduces mixed classical-quantum code patterns that require deeper contextual understanding.
-To bridge this gap, we propose the first PennyLane-centric dataset—PennyLang—paired with an advanced evaluation pipeline that incorporates RAG-based retrieval.
-II-C Retrieval-Augmented Generation and GraphRAG
-Retrieval-Augmented Generation (RAG) [47] has emerged as a popular method to ground LLM outputs using external context.
-In quantum programming, RAG helps models compensate for limited training data and outdated knowledge by retrieving relevant examples at inference time.
-However, baseline RAG systems suffer from semantic drift, irrelevant context injection, and weak code relevance due to shallow vector-based similarity matching.
-To address this, we employ GraphRAG[32], a graph-enhanced retrieval framework designed to capture structural and functional relationships between quantum code samples.
-GraphRAG models PennyLang as a graph of interconnected examples, allowing retrievals that are topologically and semantically aligned with the user query.
-Compared to traditional RAG, GraphRAG significantly improves context precision and reduces hallucinations in generated code.
-This leads to large gains in generation accuracy on the PennyLang test dataset.
+In comparison, quantum programming has seen limited adoption of LLMs, with most efforts focused on the Qiskit framework.
+IBM’s Qiskit Code Assistant [22] supports Qiskit programs’ debugging and optimizations, while [45] adapted BERT-based models for enhanced quantum code comprehension.
+CodeLlama-Quantum [27] extended Meta’s CodeLlama to address Qiskit-specific use cases.
+Despite these advances, key limitations persist.
+The field lacks large, high-quality datasets for training and evaluation, and quantum circuits demand deeper reasoning due to entanglement, measurement, and complex control logic.
+Furthermore, hybrid programming paradigms like PennyLane interleave classical and quantum instructions, requiring models to understand multi-modal code patterns and contextual dependencies.
+To address these challenges, we propose PennyLang, the first PennyLane-centric dataset, supported by an evaluation pipeline that integrates RAG to improve code relevance and contextual grounding.
+Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
 Figure 2: Methodology for PennyLang Quantum Code Generation and Evaluation.
-Data is collected from GitHub, quantum books, and PennyLane docs, then refined, tokenized, and formatted for LLMs.
-The curated dataset is processed by GPT-4o for input-output conversion.
-User queries are handled via a retrieval-augmented generation (RAG) framework using LangChain and Chroma to fetch relevant code.
-Responses from GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B-Instruct-Turbo are compared against non-RAG outputs and evaluated on functionality, syntax, and modularity to assess RAG’s impact.
-Our contributions advance AI-assisted quantum programming by enabling LLMs to provide high-quality, domain-specific assistance in PennyLane.
-TABLE II: Comparison of Related Work and Our Contributions
-Research Area Existing Works Our Contributions.
-LLMs for Code Generation Codex [26], CodeBERT [43] PennyLane-specific dataset
-LLMs for Quantum Programming Qiskit Assistant [22], Qiuchi et al. [46] Hybrid quantum-classical dataset
-Quantum Code Datasets QCircuitNet [38], QDataset  [38] First PennyLane dataset
+We collect and refine quantum code from GitHub, textbooks, and PennyLane documentation, then use GPT-4o to convert cleaned snippets into instruction–query pairs (with code and tests).
+These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts.
+GPT-4o-mini, Claude 3.5 Sonnet, LLaMa 4-17B Maverick, and Qwen 2.5-7B-Instruct-Turbo are evaluated on functional correctness to quantify the benefit of PennyLang dataset.
 III PennyLang Framework
@@
   • Forked repositories were systematically excluded.
-  • Only .py files (Python scripts) were extracted for further analysis.
+  • Only Python (.py) files were extracted for further analysis.
   • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
@@
 Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code.
-This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [40].
+This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations.
+In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [39].
 III-A2 Extracting Quantum Code from Books
-In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books:
-  • “Quantum Machine Learning: An Applied Approach” [48]
-  • “A Practical Guide to Quantum Machine Learning and Quantum Optimization” [49]
-These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs). The extraction process involved:
+In addition to GitHub, we extracted PennyLane-related code snippets from two quantum computing books, Quantum Machine Learning: An Applied Approach  [46] and A Practical Guide to Quantum Machine Learning and Quantum Optimization [47].
+These books provide theoretical foundations and applied examples of PennyLane-based quantum machine learning (QML) and variational quantum algorithms (VQAs).
+The extraction process involved:
   1. 1.
@@
   4. 4.
-    Manually verification of each snippet with explanations from the books, converting descriptive text into structured comments.
+    Manually verification with explanations from the books, converting descriptive text into structured comments.
 Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code).
@@
 III-A3 Scraping Code from PennyLane Documentation
-The official PennyLane documentation [39] serves as a comprehensive reference for quantum programming, covering:
-  • Quantum circuits and variational algorithms.
-  • Quantum operators and templates.
-  • Hybrid quantum-classical optimization.
-  • Quantum chemistry applications.
-TABLE III: Tokenized Input and Attention Masking.
-Tokenized Input Attention Mask
-"<pad>" 0
-"<pad>" 0
-"def" 1
-":" 1
-"circuit" 1
-Tokenized Input Attention Mask
-"(" 1
-")" 1
-"return" 1
-"qml.expval" 1
-To further enhance our dataset, we systematically scraped all code snippets from the official PennyLane website. This included examples embedded in tutorials and API references.
-As with book sources, documentation explanations were manually converted into comments, ensuring that each code sample contained self-contained, contextualized information.
-This practice enhances LLM interpretability, as models trained on such data can generate both code and explanations.
+The official PennyLane documentation [38] is a comprehensive resource covering topics such as quantum circuits, variational algorithms, operators and templates, hybrid quantum–classical optimization, and quantum chemistry applications.
+To enrich our dataset, we systematically extracted all code snippets from the official website, including those embedded in tutorials and API references.
+Documentation text accompanying each example was manually converted into inline comments, ensuring every sample is self-contained and context-rich.
+This approach improves LLM interpretability, enabling models trained on the dataset to produce both functional code and explanatory text.
 III-A4 Dataset Summary and Statistics
-The final dataset contains a total of 3,347 PennyLane code samples, as shown in Table IV.
-TABLE IV: Dataset Composition.
-Source Samples Collected
-Unofficial GitHub repositories 1,952
-GitHub (PennyLaneAI repository + others) 1,321
-Official PennyLane Documentation 53
-Quantum Computing Books 21
-Total Dataset Size 3,347
-Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for fine-tuning large language models (LLMs) for PennyLane-based quantum programming.
+The final dataset containing 3,347 PennyLane code samples is shown in Table III.
+TABLE III: Dataset Composition
+Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository + others) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
+Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for enhancing LLMs for PennyLane-based quantum programming.
 III-A5 Ensuring Data Integrity and Diversity
-To maintain dataset integrity, we applied multiple quality control measures:
+To keep dataset integrity, we applied multiple quality control measures:
   1. 1.
@@
 The collected dataset significantly enhances AI-driven PennyLane code assistance in multiple ways:
-  • Improving Model Performance: High-quality annotations ensure that models understand both syntax and context.
-  • Facilitating Explainable AI: Models trained on our dataset can generate explanatory comments alongside code.
+  • Improving Model Performance: High-quality annotations ensure models understand both syntax and context.
+  • Facilitating Explainable AI: Models trained on PennyLang can generate explanatory comments alongside code.
 This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
-TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework.
-Component Function Description
-DirectoryLoader Document Loader Loads PennyLane dataset samples as text files from a directory.
-Chroma Vector Database Stores embeddings for retrieval using similarity search.
-OpenAIEmbeddings Embedding Model Converts text into vector embeddings for indexing and retrieval.
-ChatPromptTemplate Prompt Formatting Defines structured prompts for AI models to ensure consistency.
-create_retrieval_chain Retrieval Pipeline Creates a pipeline that integrates stored embeddings for improved responses.
-ChatOpenAI OpenAI LLM Interface Connects to OpenAI models (GPT-4o-mini) for response generation.
-ChatAnthropic Anthropic LLM Interface Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses.
-TABLE VI: LLMs Used in the RAG-Based Evaluation Framework.
-Model Parameters Architecture Provider Usage in Pipeline
-Qwen2.5-7B-Instruct-Turbo 7B Causal Decoder-Only Transformer Alibaba Cloud Generates PennyLane code
-GPT-4o Mini Not disclosed Causal Decoder-Only Transformer OpenAI Evaluates and scores generated responses
-Claude 3.5 Sonnet Not disclosed Causal Decoder-Only Transformer Anthropic Alternative model for response generation
 III-B Tokenization and Formatting for LLM Training
@@
 III-B1 Handling Variable-Length Code Samples
-The collected PennyLane code samples exhibit varying lengths after tokenization.
+The collected PennyLane code exhibit varying lengths after tokenization.
 Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$.
@@
   • Manually Defined Padding: If a tokenizer lacks an explicit padding token, a placeholder token (e.g., "0" or "<mask>") must be manually assigned.
-  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [52], ensuring that padding tokens do not interfere with causal attention mechanisms.
-In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, and Claude 3.5 Sonnet.
-This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining autoregressive consistency.
+  • Left vs. Right Padding: Left padding is commonly used in autoregressive models [50], ensuring that padding tokens do not interfere with causal attention.
+In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, Claude 3.5 Sonnet, and LLaMa 4.
+This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
 III-B3 Applying Attention Masking
@@
 This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data.
-Table III represents the attention mask matrix applied to a batch of tokenized samples.
+Table IV shows the attention mask matrix applied to a batch of tokenized samples.
 Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
+TABLE IV: Tokenized Input and Attention Masking.
+Token Mask Token Mask Token Mask "<pad>" 0 "def" 1 "circuit" 1 "<pad>" 0 ":" 1 "(" 1 ")" 1 "return" 1 "qml.expval" 1
   • Attention Mask Structure: A binary mask is assigned to each token:
@@
   • Impact on Training: The self-attention mechanism [50] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
-Figure 3: Distribution of instruction and response lengths with their correlation.
-The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
 III-C Methodology for Dataset Evaluation
 III-C1 Retrieval-Augmented Generation (RAG) Framework
-To assess the impact of context-aware prompting on LLM-based quantum code generation, we adopt a Retrieval-Augmented Generation (RAG) framework.
-This method enhances generation quality by incorporating relevant external context retrieved from our curated PennyLang dataset.
-RAG helps address common shortcomings in LLMs, such as hallucinations and outdated or incorrect syntax, by anchoring their output in concrete examples from real-world PennyLane programs.
-This results in improved syntactic validity, functional correctness, and alignment with PennyLane best practices.
-III-C2 RAG Implementation using LangChain
-Our RAG pipeline is implemented using the LangChain framework, which provides a modular infrastructure for document loading, vector embedding, retrieval, and prompt formatting.
-The dataset—formatted in an instruction-response structure—is first embedded using the OpenAI Embeddings module.
-The resulting vectors are stored in a Chroma database for efficient retrieval.
-When a user query is submitted (e.g., a quantum programming task or challenge description), we apply Maximum Marginal Relevance (MMR) search to extract the top relevant examples from the dataset.
-These examples are concatenated into a structured prompt and fed into the LLM for code generation.
-To ensure fairness and evaluate RAG’s effectiveness, we generate responses both with and without retrieved context using the same query.
-III-C3 RAG Evaluation
-To benchmark the performance of RAG-augmented LLMs, we evaluate three state-of-the-art models: GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5-7B-Instruct-Turbo.
-Each model is tested on 29 quantum code generation tasks sampled from the QHack challenge dataset.
-For each query, both a baseline (non-RAG) and a RAG-enhanced response is generated.
-We then use GPT-4o to evaluate the outputs according to a rubric comprising three criteria:
-  • Functionality: Does the code produce the correct quantum behavior?
-  • Syntax: Is the code valid and executable in PennyLane?
-  • Modularity: Is the solution well-organized and reusable?
-Figure 4: Performance of GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5 with and without RAG. Mean scores (±1 SD) over 29 trials are shown for each model. Blue: RAG; Orange: Non-RAG.
-Figure 4 summarizes the outcome of these evaluations.
-RAG-enhanced prompting consistently improves model performance across all criteria.
-On average, GPT-4o Mini improves by 11.67%, Claude 3.5 by 7.69%, and Qwen 2.5 by 14.38%.
-These gains highlight how retrieval provides critical context, especially for complex or multi-step quantum circuits.
-III-C4 GraphRAG for Advanced Retrieval
-To further improve retrieval precision beyond vector-based similarity, we extend our RAG framework using GraphRAG.
-While the baseline RAG implementation relies solely on Maximum Marginal Relevance (MMR) vector retrieval, GraphRAG incorporates structural relationships between code samples to guide context selection.
-In this method, the PennyLang dataset is organized into a graph where nodes represent code samples and edges represent functional or structural similarity—such as shared quantum gates, observables, and template usage patterns.
-When a query is received, GraphRAG identifies a subgraph around the most semantically similar samples and performs hybrid retrieval by combining graph neighborhood traversal with traditional vector similarity.
-This enables the selection of more functionally aligned and topologically relevant code snippets, resulting in higher quality context for LLMs.
-We evaluate GraphRAG using the same task set as our baseline experiments but restrict the evaluation to the GPT-4o Mini model due to its superior baseline performance.
-The model was tested on 146 PennyLang test tasks in both no-RAG and GraphRAG settings.
-The correctness of the generated outputs was determined by whether the solution passed the associated unit tests.
-TABLE VII: Performance of GPT-4o Mini: No-RAG vs. GraphRAG (146 tasks).
-Metric No-RAG GraphRAG
-Correct Solutions 30 85
-Incorrect Solutions 116 61
-Total Samples 146 146
-Accuracy (%) 20.55% 58.22%
-As shown in Table VII, the GraphRAG approach yields a dramatic improvement in accuracy—from 20.55% without retrieval to 58.22% with graph-enhanced retrieval.
-This threefold increase highlights the value of graph-based structural reasoning in quantum programming tasks, particularly for problems involving complex circuit construction or uncommon PennyLane APIs.
-Together, these results establish GraphRAG as a significant enhancement to conventional RAG pipelines and demonstrate its practical utility in improving the performance of LLMs on challenging real-world quantum programming benchmarks.
+To improve PennyLane code generation by reducing hallucinations and outdated syntax—and to better align outputs with PennyLane best practices—we adopt a retrieval-augmented generation (RAG) [52] approach.
+We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction.
+The PennyLang dataset, formatted as instruction–response pairs, is embedded using OpenAIEmbeddings and stored in a Chroma vector database for efficient similarity search.
+At inference, user queries are embedded, relevant context is retrieved, and the combined prompt is sent to the target LLM.
+Table V summarizes the key PennyLang RAG framework components, including their role in dataset loading, embedding storage, prompt formatting, retrieval chaining, and interfacing with multiple LLM backends.
+TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
+PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o-mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
+When a user query is submitted (e.g., a quantum programming task or challenge description), we use Maximum Marginal Relevance (MMR) search to retrieve the most relevant examples from the dataset.
+The retrieved examples are concatenated into a structured prompt and passed to the LLM for code generation.
+For a fair assessment of RAG’s impact, we generate outputs for each query both with and without retrieved context under identical conditions.
 IV Statistical Evaluation of the Dataset
@@
 Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions.
+Figure 4: Distribution of instruction and response lengths with their correlation.
+The left plot shows the normal distribution of instruction lengths, the middle plot highlights the right-skewed distribution of response lengths, and the right scatter plot illustrates the relationship between instruction and response length.
 IV-B Instruction and Response Length Distributions
 Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset.
-Instruction lengths, as shown in the left plot, follow a normal distribution with a peak around 35 words.
+Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words.
 The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare.
@@
 Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs.
-The instruction-response correlation, illustrated in the rightmost scatter plot, reveals a positive correlation between instruction and response length.
+The instruction-response correlation, illustrated in the rightmost plot, reveals a positive correlation between instruction and response length.
 While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
@@
 Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
-Figure 6: Feature Group Usage Across Domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories, with color intensity reflecting count.
+Figure 5: Feature group usage across domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories (QML, Quantum Hardware, Quantum Chemistry and Optimization), with color intensity reflecting count.
+TABLE VI: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o-mini, and Claude 3.5 Sonnet on PennyLane
+Metric Qwen 7B LLaMa 4 GPT-4o-mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Total Notebooks 1320 1320 1320 1320 Working Notebooks 467 495 480 95 901 875 895 799 712 780 755 625 899 911 856 1001 Success Rate (%) 41.66 45.07 43.18 8.71 84.84 84.84 82.17 78.78 80.30 84.84 81.81 85.60 89.01 92.44 87.50 95.07 Pass@1 87 100 88 19 179 171 179 150 144 158 149 134 187 182 170 194 Pass@2 99 109 103 21 198 202 199 178 171 189 175 175 210 218 203 230 Pass@3 104 113 109 21 212 214 209 191 190 207 195 199 223 232 214 238 Pass@4 109 117 110 22 216 217 212 202 202 219 209 215 230 242 230 246 Pass@5 110 119 114 23 224 224 217 208 212 224 216 226 235 244 231 251
 IV-D Feature Category Association with Quantum Functions
@@
 IV-E Insights and Implications
-The findings from Figures 3, 5, 7, and 6 offer several key insights:
+Findings from Figures 4, 3, 6, and 5 offer several insights:
   • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
@@
   • Domain-Specific Usage: Different quantum fields rely on distinct feature sets, highlighting potential customization opportunities for specialized LLM training.
-V Case Studies
-The case studies compare base LLM-generated PennyLane code with RAG-enhanced implementations, showcasing notable improvements in functionality, syntax, and modularity across different quantum programming tasks.
-In the first task, implementing a 1-qubit parameterized RX gate, the base model introduced redundant imports, suboptimal measurement choices, and missing hardware-specific optimizations.
-The RAG-optimized implementation addressed these issues by incorporating hardware-aware device initialization, optimized measurement practices, and type annotations, achieving a perfect 10/10 functionality score.
-The second task, developing an RX gate circuit with optional parameter binding, revealed static parameter handling and the absence of deferred execution in the base LLM output.
-The RAG-enhanced version introduced flexible parameter binding, probabilistic measurement strategies, and deferred execution patterns, improving adaptability and modularity.
-The third case study, constructing a variational quantum classifier using StronglyEntanglingLayers, further demonstrated the superiority of RAG-based generation.
-The base LLM solution lacked adaptive training integration, efficient optimization, and TensorFlow-Keras support, requiring manual gradient updates.
-In contrast, the RAG-optimized model seamlessly integrated TensorFlow, Keras, and the Adam optimizer, significantly enhancing training efficiency and parameter optimization.
-Overall, across all tasks, RAG-assisted code generation consistently outperformed the base LLM, producing more structured, modular, and functionally superior PennyLane code.
-This highlights the effectiveness of retrieval-augmented approaches in enhancing quantum programming workflows.
-Task: Implement 1-qubit parameterized RX gate using PennyLane
-Base LLM Implementation (Without RAG) Limitations Identified: Unnecessary numpy import Suboptimal measurement choice (expectation value vs state) Missing hardware-specific optimizations 1import pennylane as qml 2from pennylane import numpy as np # Redundant import 3 4dev = qml.device(’default.qubit’, wires=1) 5 6@qml.qnode(dev) 7def circuit(theta): 8 qml.RX(theta, wires=0) 9 return qml.expval(qml.PauliZ(0)) # Non-optimal measurement Listing 1: Non-RAG Implementation Evaluation Results: Functionality: 9/10 (Valid but limited measurement) Syntax: 9/10 (Redundant imports) Modularity: 8/10 (Missing type hints)
-RAG-Enhanced Implementation Retrieved Context: Hardware-aware device initialization (sample_1137.txt) State measurement best practices (sample_214.txt) Testing methodologies (sample_23.txt) 1import pennylane as qml 2 3dev = qml.device("lightning.qubit", wires=1) # Hardware-optimized 4 5@qml.qnode(dev) 6def circuit(theta: float): # Type hinting 7 qml.RX(theta, wires=0) 8 return qml.state() # Full state measurement Listing 2: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Complete state analysis) Syntax: 10/10 (Clean imports) Modularity: 9/10 (Type annotations)
-Task: Create 1-qubit RX gate circuit with optional parameter binding
-Base LLM Implementation (Without RAG) Limitations Identified: Static parameter handling Missing deferred parameterization Limited measurement flexibility 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 if theta: # Binary parameter check 6 qml.RX(theta, wires=0) 7 return qml.expval(qml.PauliZ(0)) # Fixed measurement Listing 3: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (No variable binding) Syntax: 9/10 (Valid but rigid) Modularity: 8/10 (No late binding support)
-RAG-Enhanced Implementation Retrieved Context: Variable parameter techniques (sample_1137.txt) Probabilistic measurement strategies (sample_214.txt) Deferred execution patterns (sample_23.txt) 1def create_circuit(theta=None): 2 dev = qml.device("default.qubit", wires=1) 3 @qml.qnode(dev) 4 def circuit(): 5 param = theta if theta else qml.variable(’theta’) 6 qml.RX(param, wires=0) 7 return qml.probs(wires=0) # Full probability measurement Listing 4: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Flexible parameter binding) Syntax: 9/10 (Clear variable handling) Modularity: 9/10 (Supports late binding)
-Task: Implement a Variational Quantum Classifier with Strongly Entangling Layers
-Base LLM Implementation (Without RAG) Limitations Identified: Uses a static quantum circuit with fixed parameters, lacking adaptive training integration.
-Lacks integration with TensorFlow/Keras, requiring a manual optimization loop.
-Uses Gradient Descent Optimizer instead of the Adam optimizer, which may slow convergence.
-1import pennylane as qml 2import numpy as np 3from sklearn.datasets import make_classification 4from sklearn.model_selection import train_test_split 5from sklearn.preprocessing import StandardScaler 6from sklearn.metrics import accuracy_score 7 8# Generate a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Scale the features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define the quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Define the quantum circuit 21@qml.qnode(dev) 22def circuit(params, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.StronglyEntanglingLayers(params, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 5: Non-RAG Implementation Evaluation Results: Functionality: 8/10 (Works but lacks adaptive training integration) Syntax: 9/10 (Code is valid but manual optimization is not ideal) Modularity: 8/10 (Training and inference should be better structured)
-RAG-Enhanced Implementation Retrieved Context: TensorFlow-Keras integration for efficient optimization.
-Adam optimizer instead of manual updates for improved convergence.
-Improved quantum circuit design based on best practices.
-1import pennylane as qml 2import numpy as np 3import tensorflow as tf 4from sklearn.datasets import make_classification 5from sklearn.model_selection import train_test_split 6from sklearn.preprocessing import StandardScaler 7 8# Create a binary classification dataset 9X, y = make_classification(n_samples=1000, n_features=6, n_classes=2, random_state=42) 10X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42) 11 12# Standardize features 13scaler = StandardScaler() 14X_train = scaler.fit_transform(X_train) 15X_test = scaler.transform(X_test) 16 17# Define quantum device 18dev = qml.device("default.qubit", wires=6) 19 20# Quantum circuit with Strongly Entangling Layers 21@qml.qnode(dev) 22def quantum_circuit(weights, x): 23 for i in range(6): 24 qml.RX(x[i], wires=i) 25 26 qml.templates.StronglyEntanglingLayers(weights, wires=range(6)) 27 28 return qml.expval(qml.PauliZ(0)) Listing 6: RAG-Optimized Code Evaluation Results: Functionality: 10/10 (Seamless Keras integration for training) Syntax: 10/10 (Concise and efficient TensorFlow pipeline) Modularity: 9/10 (Well-structured training process)
+V Evaluation
+To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts.
+Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs.
+The evaluation utilized the Pass@5 metric, where each model generates five candidate solutions per test, and success rate (%) is computed as $\text{pass@5}/{264}\times 100$, where pass@5 indicates the number of test cases for which at least one correct solution was generated among the five attempts.
+In total, 1320 notebooks were generated and executed per model (264 tests × 5 completions), ensuring a rigorous, code-level assessment of correctness.
+Table VII lists the details for evaluated models.
+TABLE VII: LLMs Used in the RAG-Based Evaluation Framework
+Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)    $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)    $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Code generation (planned evaluation) GPT-4o Mini (OpenAI)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Alternative model for response generation
+Open-Source Models.
+Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4.
+These models showed clear improvements when used with our dataset.
+Notably, RAG with 75% context coverage often outperformed the full-context setting.
+This suggests that a moderate amount of relevant context may reduce noise and improve grounding, while excessive retrieval might introduce redundancy or irrelevant tokens.
+Commercial Models.
+For comparison, we also tested GPT-4o-mini and Claude 3.5.
+As expected, these commercial models demonstrated high performance even without RAG.
+However, the added context from our dataset provided limited benefit.
+This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.
+Discussion.
+The evaluation reveals important interactions between dataset design, model capacity, and retrieval strategies.
+Open-source models such as Qwen 7B and LLaMa 4 showed notable improvements when augmented with PennyLang, confirming that domain-specific RAG can effectively compensate for limited pretraining in quantum programming.
+Interestingly, the 75% context setting consistently outperformed the full-context retrieval, indicating that focused and moderately sized context windows improve model grounding by avoiding redundancy and context saturation.
+In contrast, commercial models like GPT-4o Mini and Claude 3.5 achieved high performance across all settings, with only marginal benefits from added retrieval.
+This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides.
+Consequently, the impact of external retrieval is diminished and can even introduce noise if not precisely targeted.
+These findings emphasize the importance of retrieval quality over quantity in RAG-based code generation.
+The consistent advantage of partial over full context suggests that carefully curated, relevant examples are more beneficial than exhaustive inclusion.
+Finally, by evaluating 1,320 generated notebooks per model, we ensure that our conclusions are grounded in executable and functionally correct outputs, a crucial criterion in the domain of quantum programming.
+These results underscore the utility of PennyLang in enhancing open-source LLM performance while providing insights into optimal RAG configurations for technical code generation tasks.
 VI Conclusion
-In this paper, we introduce PennyLang, a high-quality, instruction-response dataset tailored for PennyLane-based quantum code generation.
-We propose a robust data collection and annotation methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples.
-To assess the practical utility of the dataset, we develop a Retrieval-Augmented Generation (RAG) evaluation framework and benchmark three state-of-the-art LLMs—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B—under both RAG and non-RAG conditions.
-Our findings demonstrate consistent performance gains with RAG, especially in functionality and syntax correctness.
-Building on this, we introduce a GraphRAG extension that models structural relationships across code snippets using graph-based retrieval.
-When applied to GPT-4o Mini, GraphRAG yields a 2.8$\times$ improvement in accuracy compared to the non-RAG baseline, highlighting its potential for deeper semantic retrieval in quantum programming.
+In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation.
+We present our data collection and prepossessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples.
+To evaluate its practical utility, we develop a RAG evaluation framework and benchmark four state-of-the-art LLMs (GPT-4o Mini, Claude 3.5 Sonnet, Qwen2.5-7B, and LLaMa 4-17B Maverick) under both RAG and non-RAG settings.
+Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge.
+This further highlights the importance of calibrated retrieval budgets, as moderate settings (e.g., 75%) often outperform full context by balancing signal and noise.
 We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development.
-Future directions include training lightweight instruction-tuned models on this dataset, scaling GraphRAG across different models and quantum frameworks, and expanding toward cross-framework generalization between PennyLane, Qiskit, and Cirq.
+Future directions include training lightweight instruction-tuned models on this dataset, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.
 Acknowledgment

### v4

@@
-PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset
+A PennyLane-Centric Dataset to Enhance LLM-based Quantum Code Generation using RAG
 Abstract
@@
 To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories.
-Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline.
+Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source and commercial models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline.
 Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness.
 Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
-Index Terms:
-Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
 I Introduction
 Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [15, 46, 26].
-Breakthroughs such as Google’s demonstration of “quantum supremacy” [4], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
+Breakthroughs such as Google’s demonstration of “quantum supremacy” [2], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [39, 56, 23, 36, 21, 57, 24, 28, 42, 13, 22, 30, 37, 29, 20].
 Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [38, 53].
@@
     RAG-based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance.
-Evaluated GPT-4o mini, Claude 3.5, Qwen 7B, and LLaMa 4 Maverick 17B under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
-Summary of Key Results: Retrieval augmentation with PennyLang yields substantial relative gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66% ($\approx$32.95% relative increase), and LLaMa 4 improves from 78.78% to 84.84% ($\approx$7.7% relative increase).
-In contrast, stronger baselines (GPT-4o-mini and Claude 3.5 Sonnet) do not benefit from naive full-context retrieval, their best performance is achieved without retrieval (85.60% and 95.07%, respectively), with full retrieval resulting in slight degradation.
-Detailed analysis is presented in Section 5.
+Framework was evaluated on seven LLMs, including both open-source and commercial models under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
+Summary of Key Results: Retrieval augmentation with PennyLang yields substantial gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66%, and LLaMa 4 improves from 78.78% to 84.84%.
+In contrast, stronger baselines (commercial models) do not benefit significantly from naive full-context retrieval.
+For example, the best performance is achieved by Claude 3.5 Sonnet without retrieval (95.07%), with full retrieval resulting in slight degradation.
+Detailed analysis is presented in Section V.
 II Background and Related Work
@@
 While LLMs have demonstrated impressive capabilities in NLP [50] and automated code generation [25], their application in quantum software development remains underexplored.
-Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows.
+Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [12], with little attention to the PennyLane framework [7, 6, 5], despite its popularity in hybrid quantum-classical workflows.
 This section presents an in-depth review of related work in three key areas:
@@
 These benchmarks collectively emphasize the importance of well-curated, structured datasets in improving LLM-driven code generation.
-TABLE II: Comparison: Related Work & Our Contributions
-Research Area Existing Works Our Contributions LLMs for Code Generation Codex [26] CodeBERT [42] PennyLane-specific dataset LLMs for Quantum Programming Qiskit Assistant [22] [45] Hybrid quantum- classical dataset Quantum Code Datasets QCircuitNet [37] QDataset [37] First PennyLane dataset
 II-B LLMs for Quantum Programming
@@
 These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts.
-GPT-4o-mini, Claude 3.5 Sonnet, LLaMa 4-17B Maverick, and Qwen 2.5-7B-Instruct-Turbo are evaluated on functional correctness to quantify the benefit of PennyLang dataset.
+Both open-source and commercial models are evaluated on functional correctness to quantify the benefit of the PennyLang dataset.
+TABLE II: Comparison: Related Work & Our Contributions
+Research Area Existing Works Our Contributions LLMs for Code Generation Codex [41] CodeBERT [14] PennyLane-specific dataset LLMs for Quantum Programming Qiskit Assistant [12] [32] Hybrid quantum- classical dataset Quantum Code Datasets QCircuitNet [55] QDataSet [45] First PennyLane dataset
 III PennyLang Framework
@@
   • Only Python (.py) files were extracted for further analysis.
-  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
   • Large, autogenerated files (e.g., test logs, package files) were excluded.
@@
 This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations.
-In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [39].
+In total, 1,3211{,}321 samples were collected from the official PennyLaneAI GitHub repository [43].
 III-A2 Extracting Quantum Code from Books
@@
   4. 4.
-    Manually verification with explanations from the books, converting descriptive text into structured comments.
+    Manual verification with explanations from the books, converting descriptive text into structured comments.
 Unlike structured repositories, book content lacked a consistent format for where explanations were placed (before or after code).
@@
 III-A4 Dataset Summary and Statistics
-The final dataset containing 3,347 PennyLane code samples is shown in Table III.
+The final dataset containing 3,3473{,}347 PennyLane code samples is shown in Table III.
 TABLE III: Dataset Composition
-Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository + others) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
+Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
 Each sample varies in token length, depending on the tokenizer used. This dataset is now suitable for enhancing LLMs for PennyLane-based quantum programming.
@@
 This dataset serves as a foundation for future research in AI-assisted quantum programming for PennyLane.
-III-B Tokenization and Formatting for LLM Training
-This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
-III-B1 Handling Variable-Length Code Samples
-The collected PennyLane code exhibit varying lengths after tokenization.
-Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$.
-Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
-III-B2 Padding Strategies and Tokenizer Variations
+III-B Tokenization and Formatting for LLM
+This section details the preprocessing pipeline for preparing our dataset for (LLM), focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
+III-B1 Padding Strategies and Tokenizer Variations
 Padding strategies differ across tokenizers:
@@
   • Left vs. Right Padding: Left padding is commonly used in autoregressive models [52], ensuring that padding tokens do not interfere with causal attention.
-In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, Claude 3.5 Sonnet, and LLaMa 4.
+In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini.
 This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
-III-B3 Applying Attention Masking
+III-B2 Applying Attention Masking
 To improve transformer efficiency, an attention mask [4] was applied post-padding.
@@
 Table IV shows the attention mask matrix applied to a batch of tokenized samples.
-Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
+Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during processing.
 TABLE IV: Tokenized Input and Attention Masking.
@@
     • 0 → Padding tokens ignored by the transformer.
-  • Impact on Training: The self-attention mechanism [50] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
+  • Impact: Self-attention [52] ensures that padding tokens do not influence attention computations, thereby reducing unnecessary computations and preventing misleading dependencies.
 III-C Methodology for Dataset Evaluation
 III-C1 Retrieval-Augmented Generation (RAG) Framework
-To improve PennyLane code generation by reducing hallucinations and outdated syntax—and to better align outputs with PennyLane best practices—we adopt a retrieval-augmented generation (RAG) [52] approach.
+To improve PennyLane code generation by reducing hallucinations and outdated syntax, and to better align outputs with PennyLane best practices, we adopt a retrieval-augmented generation (RAG) [31] approach.
 We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction.
@@
 TABLE V: PennyLang Components Used in the RAG-based Evaluation Framework
-PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o-mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
+PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o Mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
 When a user query is submitted (e.g., a quantum programming task or challenge description), we use Maximum Marginal Relevance (MMR) search to retrieve the most relevant examples from the dataset.
@@
 IV-A Response Lengths Across Feature Groups
-Figure 3 provides a box plot representation of response lengths across different feature groups.
+Fig. 3 provides a box plot representation of response lengths across different feature groups.
   • Response lengths remain relatively stable across feature groups, suggesting uniformity in generated outputs.
@@
 IV-B Instruction and Response Length Distributions
-Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset.
+Fig.
+4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset.
 Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words.
@@
 IV-C Feature Group Usage Across Quantum Domains
-The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 5.
+The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Fig.
+5.
 Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation.
@@
 Figure 5: Feature group usage across domains: Heatmap illustrating how frequently each quantum feature group appears across the different domain categories (QML, Quantum Hardware, Quantum Chemistry and Optimization), with color intensity reflecting count.
-TABLE VI: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o-mini, and Claude 3.5 Sonnet on PennyLane
-Metric Qwen 7B LLaMa 4 GPT-4o-mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Total Notebooks 1320 1320 1320 1320 Working Notebooks 467 495 480 95 901 875 895 799 712 780 755 625 899 911 856 1001 Success Rate (%) 41.66 45.07 43.18 8.71 84.84 84.84 82.17 78.78 80.30 84.84 81.81 85.60 89.01 92.44 87.50 95.07 Pass@1 87 100 88 19 179 171 179 150 144 158 149 134 187 182 170 194 Pass@2 99 109 103 21 198 202 199 178 171 189 175 175 210 218 203 230 Pass@3 104 113 109 21 212 214 209 191 190 207 195 199 223 232 214 238 Pass@4 109 117 110 22 216 217 212 202 202 219 209 215 230 242 230 246 Pass@5 110 119 114 23 224 224 217 208 212 224 216 226 235 244 231 251
 IV-D Feature Category Association with Quantum Functions
-To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 6.
-Figure 6: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features.
-The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns).
-This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
 Quantum gates and operators exhibit the strongest associations with multiple features, particularly with commonly used functions such as qml.PauliZ, qml.RX, and qml.RY, which are integral to quantum circuit construction.
@@
 IV-E Insights and Implications
-Findings from Figures 4, 3, 6, and 5 offer several insights:
+Findings from Figures 3, 4, and 5 offer several insights:
   • Balanced Instruction Lengths: The normal distribution of instruction lengths ensures the dataset is well-structured for model training.
@@
 To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts.
-Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs.
-The evaluation utilized the Pass@5 metric, where each model generates five candidate solutions per test, and success rate (%) is computed as $\text{pass@5}/{264}\times 100$, where pass@5 indicates the number of test cases for which at least one correct solution was generated among the five attempts.
-In total, 1320 notebooks were generated and executed per model (264 tests × 5 completions), ensuring a rigorous, code-level assessment of correctness.
-Table VII lists the details for evaluated models.
-TABLE VII: LLMs Used in the RAG-Based Evaluation Framework
-Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)    $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)    $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Code generation (planned evaluation) GPT-4o Mini (OpenAI)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Alternative model for response generation
+The test cases (prompts) cover a broad spectrum of tasks, ranging from basic circuit construction to advanced hybrid algorithm design, encompassing state preparation, parameterized gate implementation, differentiation routines, and optimization workflows.
+Several tasks address realistic applications, including molecular energy estimation via the Variational Quantum Eigensolver (VQE) for systems such as $\mathrm{H_{2}}$ and $\mathrm{LiH}$, as well as hybrid quantum–classical models trained on toy datasets for classification and regression.
+An example of an advanced prompt is:
+“Write a PennyLane program that implements a Variational Quantum Eigensolver to estimate the ground-state energy of the $\mathrm{H_{2}}$ molecule using the default.qubit device.
+Define the molecular Hamiltonian, construct a parameterized ansatz circuit, and optimize its parameters using gradient-based optimization until convergence.”
+Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs under different context settings (100%, 75%, 50%, and 0%), where 100% denotes full retrieved context and 0% corresponds to no RAG.
+This setup was used to evaluate the effect of retrieval on code generation performance.
+The evaluation employed pass@1 through pass@5, where each model generated five candidate solutions per test.
+The success rate (%) was computed as $\texttt{pass@k}/264\times 100$, where pass@k denotes the number of test cases, out of 264, for which at least one correct solution was obtained within the first k attempts.
+A generated solution was counted as correct if the notebook executed successfully and satisfied the corresponding test-case checks for the given inputs.
+The evaluation focused on whether the generated code correctly implemented the required task and produced the expected valid output for that test case.
+For tasks involving learning or optimization, a solution was considered correct if the requested workflow executed properly and returned the required type of result, regardless of the final performance value.
+In total, 1,320 notebooks were generated and executed per model (264 tests $\times$ 5 completions), providing a rigorous code-level functional evaluation.
+Table VI lists the evaluated models.
+TABLE VI: LLMs Used in the RAG-based Evaluation Framework
+Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)     $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)     $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Code generation GPT-4o Mini (OpenAI)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Alternative model for response generation Gemini 2.5 Flash (Google DeepMind)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Rapid retrieval-augmented generation and code reasoning Claude Haiku (Anthropic)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Lightweight LLM for fast inference and evaluation tasks GPT-5 Mini (OpenAI)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Enhanced multi-turn reasoning and tool-augmented evaluation
+TABLE VII: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o Mini, and Claude 3.5 Sonnet on PennyLane.
+Results for open-source models (Qwen and LLaMa) are highlighted in purple to distinguish their RAG-based performance gains from commercial models.
+Metric Qwen 2.5 7B LLaMa 4 Maverick GPT-4o Mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Pass@1 33.0% 37.9% 33.3% 7.2% 67.8% 64.8% 67.8% 56.8% 54.5% 59.8% 56.4% 50.8% 70.8% 68.9% 64.4% 73.5% Pass@2 37.5% 41.3% 39.0% 8.0% 75.0% 76.5% 75.4% 67.4% 64.8% 71.6% 66.3% 66.3% 79.5% 82.6% 76.9% 87.1% Pass@3 39.4% 42.8% 41.3% 8.0% 80.3% 81.1% 79.2% 72.3% 72.0% 78.4% 73.9% 75.4% 84.5% 87.9% 81.1% 90.2% Pass@4 41.3% 44.3% 41.7% 8.3% 81.8% 82.2% 80.3% 76.5% 76.5% 82.9% 79.2% 81.4% 87.1% 91.7% 87.1% 93.2% Pass@5 41.7% 45.1% 43.2% 8.7% 84.8% 84.8% 82.2% 78.8% 80.3% 84.8% 81.8% 85.6% 89.0% 92.4% 87.5% 95.1%
+Figure 6: Ablation Study: Vanilla RAG vs. Graph RAG evaluated on Gemini 2.5 Flash and Claude Haiku.
+Analysis: Vanilla RAG vs. GraphRAG.
+Figure 6 compares vanilla retrieval-augmented generation (RAG) against a GraphRAG variant on our PennyLane benchmark, using two representative commercial models (Gemini 2.5 Flash and Claude Haiku).
+Across all pass@k metrics, vanilla RAG consistently outperforms GraphRAG for both models, indicating that the added graph-structuring step does not translate into better executable-code success in this setting.
+Overall, these results suggest that, for our dataset and tasks, the primary gains come from retrieving a small set of directly relevant PennyLane exemplars rather than from explicitly modeling inter-document relations via a graph.
+Since our benchmark emphasizes functional correctness of generated notebooks, where small syntactic or API-level deviations can cause execution failures, the precision of the retrieved snippets is often more important than broader relational coverage.
+Graph construction can also impose additional routing/selection constraints (e.g., traversing nodes that are only indirectly related), which may reduce the density of immediately usable code patterns in the final prompt compared to standard similarity-based retrieval.
+Given that vanilla RAG consistently outperforms GraphRAG in this ablation study, we adopt vanilla RAG as the default retrieval strategy for the remainder of the evaluations.
 Open-Source Models.
 Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4.
-These models showed clear improvements when used with our dataset.
+As shown in Table VII, these models showed clear improvements when used with our dataset.
 Notably, RAG with 75% context coverage often outperformed the full-context setting.
@@
 Commercial Models.
-For comparison, we also tested GPT-4o-mini and Claude 3.5.
-As expected, these commercial models demonstrated high performance even without RAG.
+For comparison, we also tested GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini.
+As shown in Tables VII and VIII, these commercial models demonstrated high performance even without RAG.
 However, the added context from our dataset provided limited benefit.
 This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.
+TABLE VIII: Evaluation of Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini on PennyLane under two RAG coverage settings (75% and 0%).
+Metric Gemini 2.5 Flash Claude Haiku GPT-5 Mini 75% 0% 75% 0% 75% 0% Pass@1 53.0% 55.3% 72.0% 67.4% 73.9% 74.6% Pass@2 69.3% 64.8% 81.8% 78.8% 85.6% 85.2% Pass@3 73.5% 68.9% 85.6% 84.1% 89.8% 88.3% Pass@4 79.2% 72.0% 88.3% 87.9% 90.5% 89.8% Pass@5 81.4% 75.0% 89.0% 89.8% 91.3% 92.0%
 Discussion.
@@
 Interestingly, the 75% context setting consistently outperformed the full-context retrieval, indicating that focused and moderately sized context windows improve model grounding by avoiding redundancy and context saturation.
-In contrast, commercial models like GPT-4o Mini and Claude 3.5 achieved high performance across all settings, with only marginal benefits from added retrieval.
+In contrast, commercial models like GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini achieved high performance across all settings, with only marginal benefits from added retrieval.
 This behavior is likely due to their extensive pretraining on diverse codebases, which may already include material similar to what our dataset provides.
@@
 In this paper, we introduce PennyLang, a high-quality instruction–response dataset tailored for PennyLane-based quantum code generation.
-We present our data collection and prepossessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples.
-To evaluate its practical utility, we develop a RAG evaluation framework and benchmark four state-of-the-art LLMs (GPT-4o Mini, Claude 3.5 Sonnet, Qwen2.5-7B, and LLaMa 4-17B Maverick) under both RAG and non-RAG settings.
+We present our data collection and preprocessing methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples.
+To evaluate its practical utility, we develop a RAG evaluation framework and benchmark seven state-of-the-art LLMs (both open-source and commercial models) under both RAG and non-RAG settings.
 Our findings show that PennyLang, when paired with RAG, substantially boosts performance for open-source models, while proprietary models exhibit minimal or negative gains with naive full-context retrieval, likely because their pre-training already covers overlapping PennyLane-related knowledge.
@@
 We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development.
-Future directions include training lightweight instruction-tuned models on this dataset, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.
+Future directions include training lightweight instruction-tuned models on this dataset, improving the filtering pipeline for continuously collecting high-quality new code snippets into the database, optimizing retrieval strategies across different model types, and extending to cross-framework generalization between PennyLane, Qiskit, and Cirq.
+Data Availability
+The dataset proposed in this paper is publicly available on GitHub at https://github.com/eBrain4Everyone/Pennylang.
 Acknowledgment